### v4

# High-dimensional Gaussian and bootstrap approximations for robust means

Abstract
Recent years have witnessed much progress on Gaussian and bootstrap approximations to the distribution of sums of independent random vectors with dimension $d$ large relative to the sample size $n$. However, for any number of moments $m>2$ that the summands may possess, there exist distributions such that these approximations break down if $d$ grows faster than the polynomial barrier $n^{\frac{m}{2}-1}$. In this paper, we establish Gaussian and bootstrap approximations to the distributions of winsorized and trimmed means that allow $d$ to grow at an exponential rate in $n$ as long as $m>2$ moments exist. The approximations remain valid under some amount of adversarial contamination. Our implementations of the winsorized and trimmed means do not require knowledge of $m$. As a consequence, the performance of the approximation guarantees “adapts” to $m$.
1 Introduction
Let $X_{1},\ldots,X_{n}$ be a sample of i.i.d. random vectors in $\mathbb{R}^{d}$ with mean vector $\mu$ and covariance matrix $\Sigma$. Furthermore, let $S_{n}=n^{-1/2}\sum_{i=1}^{n}\mathinner{(X_{i}-\mu)}$. Since the seminal paper of Chernozhukov et al. (2013) there has been much renewed interest in Gaussian approximations to the distribution of $S_{n}$ when $d=d(n)$ is large relative to $n$. Letting $Z\sim\mathsf{N}_{d}(0,\Sigma)$ and $\mathcal{H}$ be the class of (generalized) hyperrectangles in $\mathbb{R}^{d}$, that is the class of all sets of the form
$H=\mathinner{\bigl\{x\in\mathbb{R}^{d}\mathrel{\mathop{\ordinarycolon}}a_{j}\leq x_{j}\leq b_{j}\text{ for all }j=1,\ldots,d\bigr\}},$
where $\infty\leq a_{j}\leq b_{j}\leq\infty$ for all $j=1,\ldots,d$, increasingly refined upper bounds have been established on
$\rho_{n}\mathrel{\mathop{\ordinarycolon}}=\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(S_{n}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z\in H\bigr)}\Bigr\rvert}$ (1)
and related quantities, cf., e.g., Chernozhukov et al. (2017a); Deng and Zhang (2020); Lopes et al. (2020); Kuchibhotla and Rinaldo (2020); Das and Lahiri (2021); Koike (2021); Kuchibhotla et al. (2021); Lopes (2022); Chernozhukov et al. (2022); Fang et al. (2023); Chernozhukov et al. (2023b); Koike (2024). We refer to the review in Chernozhukov et al. (2023a) for further references. For example, when the entries of $X_{i}=(X_{i,1},\ldots,X_{i,d})^{\prime}$ are (uniformly) sub-exponential, $\rho_{n}\to 0$ if $\log(d)=o\mathinner{\bigl(n^{1/5}\bigr)}$, cf. Chernozhukov et al. (2022). Thus, $d$ can grow exponentially fast with $n$ and this rate can be further improved under additional assumptions on the distribution of the $X_{i}$ such as, e.g., variance decay conditions on $\Sigma$ as in Lopes et al. (2020) or eigenvalue conditions as in Fang and Koike (2021); Kuchibhotla and Rinaldo (2020); Chernozhukov et al. (2023b).
Despite the progress on such high-dimensional Gaussian approximations for $S_{n}$, it follows from Remark 2 in Zhang and Wu (2017) and Theorem 2.1 in Kock and Preinerstorfer (2024) that for every $m\in(2,\infty)$ there exist i.i.d. random vectors $X_{1},\ldots,X_{n}$ with independent entries $X_{i,j}\sim P_{m}$, and $P_{m}$ depending neither on $n$ nor $d$, having mean zero, variance one, and finite $m$th absolute moment, such that if for some $\xi\in(0,\infty)$ it holds that $\limsup_{n\to\infty}\frac{d}{n^{m/2-1+\xi}}>0$, then $\limsup_{n\to\infty}\rho_{n}=1$. In particular, for any given $m\in(2,\infty)$, the Gaussian approximation $\rho_{n}\to 0$ does not hold uniformly over all distributions with bounded $m$th moments when $d$ grows exponentially in $n$. Figure 1 illustrates how the Gaussian approximation to $S_{n}$ can fail severely over $\mathcal{H}$ when the $X_{i,j}$ have relatively heavy tails. In particular, letting $x||_{\infty}=\max_{1\leq j\leq d}|x_{j}|$ for $x\in\mathbb{R}^{d}$, it is seen that $S_{n}||_{\infty}$ has much heavier tails than $Z||_{\infty}$.
Conversely, it is a simple consequence of Theorem 2 in Chernozhukov et al. (2023a), cf. Theorem 2.2 in Kock and Preinerstorfer (2024), that $\rho_{n}\to 0$ uniformly over a large class of distributions with bounded $m$th moments if there exists a $\xi\in(0,\infty)$ such that $\lim_{n\to\infty}\frac{d}{n^{m/2-1-\xi}}=0$. Hence, a critical phase transition occurs for the asymptotic behaviour of $\rho_{n}$ at $d=n^{m/2-1}$. As $d$ passes this threshold from below, the limit of $\rho_{n}$ jumps from zero to one.
Motivated by this phase transition, Resende (2024) recently studied the case where $S_{n}$ is replaced by a suitably trimmed mean and $\mathcal{H}$ is replaced by the subfamily $\mathcal{R}$ consisting of all sets of the form $R=\mathinner{\{x\in\mathbb{R}^{d}\mathrel{\mathop{\ordinarycolon}}x_{j}\leq t\text{ for all }j=1,\ldots,d\}}$, where $t\in\mathbb{R}$. He obtained Gaussian approximations that are informative even when the $X_{i}$ only possess $m>2$ moments and $d$ grows exponentially fast in $n$. This is of fundamental importance, as it shows that one can break through the barrier $d=n^{m/2-1}$ faced by $\rho_{n}$, which is based on $S_{n}$. The exact permitted growth rate of $d$ depends on $m$. As $m\to\infty$, his result allows $d$ to grow almost as fast as $\exp(n^{1/6})$ for the Gaussian approximation as well as an empirical bootstrap and as fast as $\exp(n^{1/8})$ for a multiplier bootstrap. Furthermore, the trimming ensures that these approximations remain valid even when some of the $X_{i}$ have been adversarially contaminated prior to being given to the statistician. This is in stark contrast to statistics based on the sample mean $S_{n}$, which have a breakdown point of $1/n$ ($S_{n}$ can be changed to any value by manipulating only one of the vectors $X_{i}$).
We also mention the work of Liu and Lopes (2024) who, motivated by the poor performance of $S_{n}$ in the presence of heavy tails, even established a dimension-independent bootstrap approximation over $\mathcal{R}$ for certain robust max-statistics related to the winsorized means we study (but working with a “variance-based” instead of a “quantile-based” winsorization) under the conditions of $L^{4}$-$L^{2}$ moment equivalence, a variance decay condition on $\Sigma$, and restrictions on the Frobenius norm of certain submatrices of the correlation matrix of the $X_{i}$. Robustness to outliers or other sources of (adversarial) data contamination were not investigated.
$n=200$ and d=5,000d=5{,}000               n=1,000n=1{,}000 and d=10,000d=10{,}000
[…]
In the present paper, we obtain Gaussian approximations to the distributions of winsorized and trimmed means over $\mathcal{H}$ without imposing any structural assumptions on $\Sigma$ (apart from positive variances). Our Gaussian approximations are valid over all of $\mathcal{H}$ rather than the smaller family $\mathcal{R}$ studied in Resende (2024), which is crucial for the study of power properties of max-type tests based on our winsorized and trimmed means; cf. the results in the companion paper Kock and Preinerstorfer (2026). In contrast to the present paper, which focuses exclusively on the canonical problem of Gaussian approximations in $\mathbb{R}^{d}$Resende (2024) also considers Gaussian approximations over VC-subgraph classes of functions and applies his results to vector mean estimation under general norms.
One limitation of the trimmed mean studied in (Theorem 2) in Resende (2024) is that the amount of trimming needed depends on $m$, which is typically unknown in practice. If one implements the trimmed mean studied in Resende (2024) based on an $m$ higher than the actual number of moments that the $X_{i}$ possess, one does not have any approximation guarantees, whereas the guarantees one obtains may be suboptimal if the $X_{i}$ possess more moments than the $m$ used in the implementation of the trimmed mean. What is more, even when $m$ is known, it is often the case that the required number of trimmed observations for the estimator studied in Resende (2024)
  • i)
    either exceeds the sample size $n$, making that trimmed mean unimplementable (cf. the left panel of Figure 1 where the trimmed mean studied in Resende (2024) is absent);
  • ii)
    or is smaller than $n$ but so large that the Gaussian approximation to the distribution of that trimmed mean is not practically useful (cf. the right panel of Figure 1 showing that the tails of the trimmed mean are much lighter than those of the approximating Gaussian due to excessive trimming).
In contrast, to implement the winsorized and trimmed means we investigate, one does not need to know the number of moments $m$ that the $X_{i}$ possess. The winsorization/trimming used in the present article draws from ideas in Lugosi and Mendelson (2021) and Kock and Preinerstorfer (2025), is quantile-based, and “adapts” to $m$. Figure 1 illustrates, in particular, that the winsorized mean we study obeys a tight Gaussian approximation even for rather large $d$.
Furthermore, our results only require $\log(d)=o\mathinner{\bigl(n^{\frac{m-2}{5m-2}}\bigr)}$, which is exponential for all $m>2$, albeit with a small exponent for $m$ close to two. As $m\to\inftydd$ is allowed to grow almost as fast as $\exp(n^{1/5})$, improving on the rate in Resende (2024), and “recovering” the best known rate for Gaussian approximations based on the sample mean of $X_{i}$ with sub-exponential entries (cf. Remark 2 in Chernozhukov et al. (2023a)). This rate remains valid for the bootstrap procedures that we consider and all results are robust to some adversarial contamination.
Our bootstrap approximations are based on a novel “winsorized” covariance matrix estimator, for which we establish performance guarantees in the presence of adversarial contamination and only requiring $m>2$ moments. Because the winsorization mechanism used in the construction of that covariance estimator is quantile-based, the estimator does not require knowledge of any unknown population quantities. In particular, its performance guarantees adapt to the unknown $m$, which may be of independent interest.
Gaussian approximations for winsorized means are obtained in Section 2, whereas bootstrap approximations are studied in 3. Some results on trimmed means are established in Section 5. Numerical results in addition to the ones already discussed in Figure 1 above are presented in Section 6.
2 Gaussian approximations for winsorized means
We first present our approximations to the distributions of high-dimensional winsorized means. Section 5 outlines the corresponding results for the version of the trimmed means we study.
Recall that $X_{1},\ldots,X_{n}$ is a sample of i.i.d. random vectors in $\mathbb{R}^{d}$, and where $X_{i}=(X_{i,1},\ldots,X_{i,d})^{\prime}$ for $i=1,\ldots,n$. Let $\mu=(\mu_{1},\ldots,\mu_{d})^{\prime}=\mathbb{E}X_{1}$, $\Sigma$ be the covariance matrix of $X_{1}$, and for $m\in[2,\infty)$ let $\sigma_{m,j}^{m}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}|X_{1,j}-\mu_{j}|^{m}$, all of which are well-defined under Assumption 2.1 below. We suppress the dependence of $d=d(n)$ on $n$ in our notation.
In this section, our main focus is to establish Gaussian approximations for winsorized means that are valid for $d$ growing exponentially in $n$ imposing only that the $X_{1,j}$ possess $m>2$ moments, $j=1,\ldots,d$. An added benefit of the winsorization is that the Gaussian approximations are robust to some amount of adversarial contamination. Under such contamination an adversary inspects the sample and returns a corrupted sample $\tilde{X}_{1},\ldots,\tilde{X}_{n}$ to the statistician satisfying that
$\mathinner{\!\bigl\lvert\mathinner{\bigl\{i\in\mathinner{\{1,\ldots,n\}}\mathrel{\mathop{\ordinarycolon}}\tilde{X}_{i}\neq X_{i}\bigr\}}\bigr\rvert}\leq\overline{\eta}_{n}n,$ (2)
where $\overline{\eta}_{n}\in(0,1/2)$ is a non-random and known upper bound on the fraction of contaminated observations. Which of $\tilde{X}_{i}$ differ from $X_{i}$ as well as their values can depend on the uncontaminated sample $X_{1},\ldots,X_{n}$. Adversarial contamination has become a popular criterion to evaluate robustness of a statistic against, as it allows for many forms of data manipulation, cf. Lai et al. (2016), Cheng et al. (2019), Diakonikolas et al. (2019), Hopkins et al. (2020), Lugosi and Mendelson (2021), Minsker and Ndaoud (2021), Bhatt et al. (2022), Depersin and Lecué (2022), Dalalyan and Minasyan (2022), Minasyan and Zhivotovskiy (2023), Minsker (2023), Oliveira et al. (2025). The recent book by Diakonikolas and Kane (2023) provides further references and discussion of various contamination settings. Since the sample mean has a breakdown point of $1/n$, Gaussian approximations based on $S_{n}$ are not robust to adversarial contamination (or large outliers).
In all asymptotic statements $n\to\infty$. Throughout, we impose the following assumption (for various values of $m$).
Assumption 2.1.
The $X_{1},\ldots,X_{n}$ are i.i.d. random vectors in $\mathbb{R}^{d}$ with $\mathbb{E}|X_{1,j}|^{m}<\infty$ for some $m\in(2,\infty)$ and all $j=1,\ldots,d$. There exist strictly positive constants $b_{1}$ and $b_{2}$ such that $\min_{j=1,\ldots,d}\sigma_{2,j}\geq b_{1}$ and $\sigma_{m}\mathrel{\mathop{\ordinarycolon}}=\max_{j=1,\ldots,d}\sigma_{m,j}\leq b_{2}$. The actually observed adversarially contaminated random vectors (in $\mathbb{R}^{d}$) are denoted $\tilde{X}_{1},\ldots,\tilde{X}_{n}$ and satisfy (2).
Imposing lower and upper bounds on moments of the $X_{1,j}$ is commonplace when establishing upper bounds on $\rho_{n}$ in (1), cf., e.g., the results in the overview Chernozhukov et al. (2023a). Let us emphasize that all of our results are valid (in particular) absent adversarial contamination, i.e., for $\overline{\eta}_{n}=0$, which is the case studied in the literature on upper bounds on $\rho_{n}$ summarized in the introduction.
2.1 The winsorized means
For real numbers $x_{1},\ldots,x_{n}$, denote by $x_{1}^{*}\leq\ldots\leq x_{n}^{*}$ their non-decreasing rearrangement. Let $\infty<\alpha\leq\beta<\infty$ and
$\phi_{\alpha,\beta}(x)=\begin{cases}\alpha\qquad\text{if }x<\alpha\\ x\qquad\text{if }x\in[\alpha,\beta]\\ \beta\qquad\text{if }x>\beta.\end{cases}$
We establish Gaussian and bootstrap approximations to the distribution of centered winsorized means $S_{n,W}\in\mathbb{R}^{d}$ where
$S_{n,W,j}=n^{-1/2}\sum_{i=1}^{n}\mathinner{\bigl(\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}(\tilde{X}_{i,j})-\mu_{j}\bigr)},\qquad j=1,\ldots,d,$ (3)
with $\hat{\alpha}_{j}=\tilde{X}_{\lceil\varepsilon_{n}n\rceil,j}^{*}$ and $\hat{\beta}_{j}=\tilde{X}_{\lceil(1-\varepsilon_{n})n\rceil,j}^{*}$ for a carefully chosen $\varepsilon_{n}\in(0,1/2)$. Thus, for each coordinate $j$, the winsorization points $\hat{\alpha}_{j}$ and $\hat{\beta}_{j}$ are order statistics of the contaminated data $\tilde{X}_{1,j},\ldots,\tilde{X}_{n,j}$. In this sense, the winsorization mechanism employed is quantile-based. [^2] For the purpose of constructing estimators of $\mu\in\mathbb{R}$ with finite-sample sub-Gaussian concentration properties, related winsorized mean estimators were recently studied in Lugosi and Mendelson (2021) and Kock and Preinerstorfer (2025). The winsorized estimators studied in the latter article remove some practical limitations of the ones studied in the former. We therefore mainly build on the results in Kock and Preinerstorfer (2025). Under adversarial contamination it is clear that even $S_{n,W}$ can perform arbitrarily badly unless at least the smallest and largest $\overline{\eta}_{n}n$ observations are winsorized. Thus, one must choose $\varepsilon_{n}\geq\overline{\eta}_{n}$. Our Gaussian approximations below apply for any $\varepsilon_{n}\in(0,0.5)$ of the form
$\varepsilon_{n}=\lambda_{1}\cdot\overline{\eta}_{n}+\lambda_{2}\cdot\frac{\log(dn)}{n},\qquad\text{with }\lambda_{1}\in(1,\infty)\text{ and }\lambda_{2}\in(0,\infty).$ (4)
We impose $\lambda_{1}>1$ since otherwise even $S_{n,W}$ can be arbitrarily manipulated by the adversary. To make efficient use of the data, we recommend $\lambda_{1}$ close to but greater than one. In the leading special case of $\overline{\eta}_{n}=0$, which is the setting in which high-dimensional Gaussian approximations for the sample mean based $S_{n}$ have been studied (cf. the literature summarized in the introduction), one has that $\varepsilon_{n}=\lambda_{2}\cdot\log(dn)/n$. The simulations in Section 6 suggest that $\lambda_{2}\approx 0.1$ works well in practice.
Without further mentioning, $\lambda_{1}\in(1,\infty)$ and $\lambda_{2}\in(0,\infty)$ are taken as fixed throughout the remainder of the paper.
2.2 Gaussian approximation
We now present a high-dimensional Gaussian approximation result for $S_{n,W}$ over $\mathcal{H}$ (defined prior to (1)) in the form of an upper bound on
$\rho_{n,W}\mathrel{\mathop{\ordinarycolon}}=\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(S_{n,W}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z\in H\bigr)}\Bigr\rvert},\qquad\text{where }Z\sim\mathsf{N}_{d}(0,\Sigma).$
We assume throughout that $d\geq 2$ and $n>3$ (such that, e.g., $\sqrt{\log(d)}>0$).
Theorem 2.1.
Let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n}\in(0,1/2)$, with $\varepsilon_{n}$ as in (4), then, for a constant $C$ depending only on $b_{1},b_{2},\lambda_{1},\lambda_{2}$, and $m$, we have
$\rho_{n,W}$ $\leq C\mathinner{\Biggl(\mathinner{\biggl[\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr]}^{\frac{1}{4}}+\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr]}\sqrt{n\log(d)}\Biggr)}$ (5)
$\quad+C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{2}{m}}\biggr]}\Biggr)}^{1/2}.$
In particular, $\rho_{n,W}\to 0$ if $\sqrt{n\log(d)}\overline{\eta}_{n}^{1-\frac{1}{m}}\to 0$ and $\log(d)/n^{\frac{m-2}{5m-2}}\to 0$.
Consider the case of $\overline{\eta}_{n}=0$. Theorem 2.1 then shows that winsorized means can break through the polynomial growth rate barrier $d=n^{m/2-1}$ that $d$ must obey for the Gaussian approximation error to the distribution of $S_{n}$, i.e., $\rho_{n}$ in (1), to converge to zero. In particular, $\rho_{n,W}\to 0$ if only $\log(d)=o\mathinner{\bigl(n^{\frac{m-2}{5m-2}}\bigr)}$, which allows for $d$ growing exponentially in $n$ for any $m>2$, albeit with a small exponent for $m$ close to two. Thus, the winsorized mean obeys a Gaussian approximation result over $\mathcal{H}$ under heavy tails and adversarial contamination, which is in analogy to the Gaussian approximation result over $\mathcal{R}\subsetneq\mathcal{H}$ in Resende (2024) for the trimmed mean analyzed there. However, let us highlight the following key differences, which we already commented on in the introduction: The implementation of the trimmed mean estimator in Resende (2024) requires knowledge of the (typically) unknown $m$. Thus, if the $X_{i}$ have fewer moments than the chosen $m$, then there are no approximation guarantees. If, on the other hand, the $X_{i}$ have more moments than the specified $m$, then the guarantees may be suboptimal. In contrast, the implementation of our winsorized mean does not depend on $m$ — it “adapts” to it. Furthermore, we recover the rate $d=o\mathinner{\bigl(\exp(n^{1/5})\bigr)}$ as $m\to\infty$, which is currently the best available for Gaussian approximations based on $S_{n}$ with sub-exponential $X_{i}$, instead of $d=o\mathinner{\bigl(\exp(n^{1/6})\bigr)}$ for the trimmed mean analyzed in Resende (2024). To make this more concrete, consider the case where the entries of the $X_{i}$ are heavy-tailed in the sense of possessing exactly, e.g., $m=4$ moments. Then, $\rho_{n,W}\to 0$ if $\log(d)=o(n^{1/9})$, whereas the trimmed mean analyzed in Theorem 2 in Resende (2024) allows $\log(d)=o(n^{3/35})$ if implemented with $m=3$, $\log(d)=o(n^{4/35})$ if implemented with the true $m=4$, and provides no guarantees if implemented with $m>4$. Thus, unless one has reliable information on the number of moments the data possesses (in this case: knows $m\approx 4$), our method adapting to $m$ allows for larger $d$. Finally, Section 6 (cf. also Figure 1 in the Introduction) shows that the trimmed mean of Resende (2024) i) often requires trimming more observations than the sample size $n$, rendering it unimplementable, or ii) its distribution is very poorly approximated by that of $Z$. In contrast, the distribution of $S_{n,W}$ is closely approximated by that of $Z$.
To prove Theorem 2.1, we first show that the order statistics $\hat{\alpha}_{j}$ and $\hat{\beta}_{j}$ can (essentially) be replaced by closely related population quantiles $Q_{\varepsilon_{n},j}$ and $Q_{1-\varepsilon_{n},j}$ of the $X_{1,j}$, such that one can analyze the non-random winsorization functions $\phi_{n,j}(x)\mathrel{\mathop{\ordinarycolon}}=\phi_{Q_{\varepsilon_{n},j},Q_{1-\varepsilon_{n},j}}(x)$ instead of $\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}(x)$ for all $j=1,\ldots,d$: [^3] A similar control of the order statistics was also used in Lugosi and Mendelson (2021) and Kock and Preinerstorfer (2025). Resende (2024) studied the trimmed mean by relating it to a winsorized mean.
$S_{n,W,j}\approx\underbrace{\frac{1}{n^{1/2}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{n,j}(\tilde{X}_{i,j})-\mu_{j}\bigr]}}_{T_{n,j}}$ $=\underbrace{\frac{1}{n^{1/2}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{n,j}(\tilde{X}_{i,j})-\phi_{n,j}(X_{i,j})\bigr]}}_{I_{n,j,1}}$
$+\underbrace{\frac{1}{n^{1/2}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{n,j}(X_{i,j})-\mathbb{E}\phi_{n,j}(X_{i,j})\bigr]}}_{I_{n,j,2}}$
$+\underbrace{\frac{1}{n^{1/2}}\sum_{i=1}^{n}\mathinner{\bigl[\mathbb{E}\phi_{n,j}(X_{i,j})-\mu_{j}\bigr]}}_{I_{n,j,3}}.$
The term $\max_{j=1,\ldots,d}|I_{n,j,1}|$ isolates the effect of the adversarial contamination of the data and $\max_{j=1,\ldots,d}|I_{n,j,3}|$ quantifies how far the winsorized means $\mathbb{E}\phi_{n,j}(X_{1,j})$ are from the population means $\mu_{j}$ of interest. Next, note that $I_{n,j,2}$ is a sample average of bounded i.i.d. random variables. Thus, letting $Z_{n}\sim\mathsf{N}_{d}(0,\Sigma_{\phi_{n}})$ with $\Sigma_{\phi_{n}}$ being the covariance matrix of $I_{n,2}=(I_{n,1,2},\ldots,I_{n,d,2})^{\prime}$, one can apply, e.g., the Gaussian approximation for sums of sub-exponential random vectors from Chernozhukov et al. (2022) to this term to show that
$\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(I_{n,2}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z_{n}\in H\bigr)}\Bigr\rvert}\qquad\text{is small}.$
Furthermore, we show that $\max_{1\leq j,k\leq d}\mathinner{\lvert\Sigma_{\phi_{n},j,k}-\Sigma_{j,k}\rvert}$ is sufficiently small for the Gaussian-to-Gaussian comparison inequality as stated in Proposition 2.1 in the previous reference (cf. also Proposition 2 in Chernozhukov et al. (2023a)) to imply that
$\mathbb{P}\mathinner{\bigl(Z_{n}\in H\bigr)}\qquad\text{in the previous display can be replaced by}\qquad\mathbb{P}\mathinner{\bigl(Z\in H\bigr)}.$
Finally, we show that for $l\in\mathinner{\{1,3\}}$ one has that $\max_{j=1,\ldots,d}|I_{n,j,l}|$ are sufficiently small for a Gaussian anti-concentration inequality to imply that these can be “ignored”. Therefore, $\mathbb{P}\mathinner{\bigl(I_{n,2}\in H\bigr)}$ can be replaced by $\mathbb{P}\mathinner{\bigl(T_{n}\in H\bigr)}$ in the penultimate display, where $T_{n}=(T_{n,1},\ldots,T_{n,d})^{\prime}$.
3 Bootstrap approximations for winsorized means
Similarly to Gaussian approximations for $S_{n}$, cf. the references in the introduction, the one for $S_{n,W}$ in Theorem 2.1 is not directly useful for statistical inference since the covariance matrix $\Sigma$ of the approximating distribution $\mathsf{N}_{d}(0,\Sigma)$ is typically unknown. Because the Gaussian-to-Gaussian comparison inequality as stated in Proposition 2.1 in Chernozhukov et al. (2022) (cf. also Proposition 2 in Chernozhukov et al. (2023a)) shows that for $Z_{1}\sim\mathsf{N}_{d}(0,\Sigma^{(1)})$ and $Z_{2}\sim\mathsf{N}_{d}(0,\Sigma^{(2)})$ with $\min_{j=1,\ldots,d}\Sigma^{(2)}_{j,j}>b$ for some $b>0$, it holds that
$\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(Z_{1}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z_{2}\in H\bigr)}\Bigr\rvert}\leq C\mathinner{\Bigl(\log^{2}(d)\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\Sigma^{(1)}_{j,k}-\Sigma^{(2)}_{j,k}\bigr\rvert}\Bigr)}^{1/2},$ (6)
for some $C=C(b)$, one can approximate the unknown $\mathbb{P}\mathinner{(Z\in H)}$ from Theorem 2.1 if an estimator $\hat{\Sigma}_{n}$ satisfying an upper bound on $\log^{2}(d)\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\hat{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}$ can be exhibited. The sample covariance matrix can be used for $d$ growing exponentially in $n$ when i) $\overline{\eta}_{n}=0$ and ii) the $X_{i}$ have sub-exponential entries. [^4] See Section 4.1 of Kuchibhotla and Chakrabortty (2022) for properties of the sample covariance matrix when the $X_{i}$ have sub-Weibull entries (generalizing sub-exponential distributions). However, since we allow for $\overline{\eta}_{n}>0$ and only impose the existence of $m>2$ moments, the sample covariance matrix cannot be used in our context.
There has been a recent interest in constructing estimators of $\Sigma$ which perform well under heavy tails (and frequently also adversarial contamination). For example, estimators with precision guarantees in the entrywise maximal distance $\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\hat{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}$ needed in (6), have been proposed in Ke et al. (2019) in the setting of heavy-tailed $X_{i}$. These estimators are based on, e.g., entrywise truncation or the median-of-means principle and the practical choice of the needed tuning parameters (which depend on unknown population quantities) is also discussed there.
In the next section, we construct an estimator $\tilde{\Sigma}_{n}$, say, which is based on suitably winsorized observations $\tilde{X}_{i}$, and for which we establish performance guarantees even for $\bar{\eta}_{n}>0$ and when the $X_{i}$ possess only $m>2$ moments. Our estimator does not depend on unknown population quantities and its performance guarantees “adapt” to the unknown $m$. We stress that Theorem 3.2 below is modular in the sense that it remains valid for any other estimator $\hat{\Sigma}_{n}$ satisfying a bound as in Theorem 3.1.
3.1 Estimating $\Sigma$
Imposing only $m>2$ moments to exist we now construct an estimator of $\Sigma$ with precision guarantees in the maximal entrywise norm. Set
$\varepsilon_{n}^{\prime}=\lambda_{1}^{\prime}\cdot\overline{\eta}_{n}+\lambda_{2}^{\prime}\cdot\frac{\log(d^{2}n)}{n},\qquad\lambda_{1}^{\prime}\in(1,\infty)\text{ and }\lambda_{2}^{\prime}\in(0,\infty).$ (7)
Like for $\lambda_{1}$ and $\lambda_{2}$, we take $\lambda_{1}^{\prime}$ and $\lambda_{2}^{\prime}$ as fixed throughout.
Writing $\hat{a}_{j}=\tilde{X}_{\lceil\varepsilon_{n}^{\prime}n\rceil,j}^{*}$$\hat{b}_{j}=\tilde{X}_{\lceil(1-\varepsilon_{n}^{\prime})n\rceil,j}^{*}$ and $\tilde{\mu}_{n,j}=n^{-1}\sum_{i=1}^{n}\phi_{\hat{a}_{j},\hat{b}_{j}}(\tilde{X}_{i,j})$ for $j=1,\ldots,d$, define $\tilde{\Sigma}_{n}$ as the matrix with entries
$\tilde{\Sigma}_{n,j,k}=\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\hat{a}_{j},\hat{b}_{j}}(\tilde{X}_{i,j})-\tilde{\mu}_{n,j}\bigr]}\mathinner{\bigl[\phi_{\hat{a}_{k},\hat{b}_{k}}(\tilde{X}_{i,k})-\tilde{\mu}_{n,k}\bigr]},\qquad 1\leq j,k\leq d.$ (8)
The estimator $\tilde{\Sigma}_{n}$, which is a winsorized analog of the sample covariance matrix, is positive semi-definite and symmetric by virtue of being a Gram matrix. Solely for the purpose of stating the precision guarantee for $\tilde{\Sigma}_{n}$ in Theorem 3.1 below, we introduce the condition
$2\varepsilon_{n}^{\prime}+\frac{\log(d^{2}n)}{n}+\sqrt{\mathinner{\Bigl(\frac{\log(d^{2}n)}{n}\Bigr)}^{2}+4\frac{\log(d^{2}n)}{n}\varepsilon_{n}^{\prime}}<1.$ (9)
We stress that (9) is not needed for any of the subsequent Gaussian and bootstrap approximations of this paper to hold, as these will be argued to be trivially true when (9) is violated, cf., e.g., the closing paragraph of the proof of Theorem 3.2. Note also that in the leading special case of $\overline{\eta}_{n}=0$, which is always imposed in the literature when studying Gaussian and bootstrap approximations to the distribution of $S_{n}$, (9) reduces to
$\mathinner{\Bigl(2\lambda_{2}^{\prime}+1+\sqrt{1+4\lambda_{2}^{\prime}}\Bigr)}\frac{\log(d^{2}n)}{n}<1,$
which is often satisfied for $\lambda_{2}^{\prime}$ small.
Remark 3.1.
The condition in (9) is actually just a conservative (simple) sufficient condition for the following milder condition: Writing $A_{+}^{\prime}=1-\lambda_{1}^{\prime-1}\mathds{1}\mathinner{\{\overline{\eta}_{n}>0\}}\in(0,1]$$A_{-}^{\prime}=1+\lambda_{1}^{\prime-1}\mathds{1}\mathinner{\{\overline{\eta}_{n}>0\}}\in[1,\infty)$, and denoting by $W_{0}$ and $W_{-1}$ the principal and lower branch of Lambert’s $W$ function (cf., e.g., Corless et al. (1996)), respectively, (9) can be replaced by
$\varepsilon_{n}^{\prime}\mathinner{\biggl(-A_{+}^{\prime}W_{0}\mathinner{\Bigl(-e^{-(\frac{\log(d^{2}n)}{\varepsilon_{n}^{\prime}n}+A_{+}^{\prime})/A_{+}^{\prime}}\Bigr)}-A_{-}^{\prime}W_{-1}\mathinner{\Bigl(-e^{-(\frac{\log(d^{2}n)}{\varepsilon_{n}^{\prime}n}+A_{-}^{\prime})/A_{-}^{\prime}}\Bigr)}\biggr)}<1.$
By (G.6) of Lemma G.6 in the appendix, the left-hand side of the previous display is upper bounded by the left-hand side of (9), explaining the condition in (9). Note that (9) also implies $\frac{\log(d^{2}n)}{n}<1$, which we repeatedly use in the proofs.
Theorem 3.1.
Let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n}^{\prime}$ is chosen as in (7) and satisfies (9), then for a constant $C$ depending only on $b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$ and $m$, we have
$\mathbb{P}\mathinner{\Biggl(\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}>C\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\Biggr)}\leq\frac{24}{n}.$ (10)
Theorem 3.1 shows that for any $m>2$ it is possible for $\max_{1\leq j,k\leq d}\mathinner{\lvert\tilde{\Sigma}_{n,j,k}-\Sigma_{j,k}\rvert}$ to converge to zero in probability even when $d$ grows exponentially in $n$. Finally, we reiterate that the approximation in Theorem 3.2 below remains valid for any covariance estimator satisfying (10).
3.2 Bootstrap consistency
Equipped with the estimator $\tilde{\Sigma}_{n}$, the following theorem justifies approximating $\mathbb{P}\mathinner{\bigl(S_{n,W}\in H\bigr)}$ for $H\in\mathcal{H}$ by sampling repeatedly from $\mathsf{N}_{d}(0,\tilde{\Sigma}_{n})$.
Theorem 3.2.
Let Assumption 2.1 be satisfied with $m>2$. Let $n>24$. If $\varepsilon_{n},\varepsilon_{n}^{\prime}\in(0,1/2)$, with $\varepsilon_{n}$ as in (4) and $\varepsilon_{n}^{\prime}$ as in (7), and $\tilde{Z}\sim\mathsf{N}_{d}(0,\tilde{\Sigma}_{n})$ conditionally on $\tilde{X}_{1},\ldots,\tilde{X}_{n}$, it holds with probability at least $1-\frac{24}{n}$ that
$\tilde{\rho}_{n,W}\mathrel{\mathop{\ordinarycolon}}$ $=\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(S_{n,W}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(\tilde{Z}\in H\mid\tilde{X}_{1},\ldots,\tilde{X}_{n}\bigr)}\Bigr\rvert}$
$\leq\mathfrak{A}_{n}+C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\Biggr)}^{1/2},$
where $\mathfrak{A}_{n}$ is the upper bound on $\rho_{n,W}$ in (5) of Theorem 2.1 and $C$ is a constant depending only on $b_{1},b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$ and $m$.
In particular, $\tilde{\rho}_{n,W}\to 0$ in probability if $\sqrt{n\log(d)}\overline{\eta}_{n}^{1-\frac{1}{m}}\to 0$ and $\log(d)/n^{\frac{m-2}{5m-2}}\to 0$.
Draws from $\mathsf{N}_{d}(0,\tilde{\Sigma}_{n})$ can be obtained efficiently by, e.g., the following multiplier bootstrap: Let $\xi_{1},\ldots,\xi_{n}$ be i.i.d. $N_{1}(0,1)$, independent of $\tilde{X}_{1},\ldots,\tilde{X}_{n}$, and set
$S_{n,\text{MB}}=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\xi_{i}V_{i}\qquad\text{where}\qquad V_{i}=\mathinner{\bigl(\phi_{\hat{a}_{1},\hat{b}_{1}}(\tilde{X}_{i,1})-\tilde{\mu}_{n,1},\ldots,\phi_{\hat{a}_{d},\hat{b}_{d}}(\tilde{X}_{i,d})-\tilde{\mu}_{n,d}\bigr)}^{\prime}\in\mathbb{R}^{d}.$
Observe that conditionally on $\tilde{X}_{1},\ldots,\tilde{X}_{n}$ the distribution of $S_{n,\text{MB}}$ is $\mathsf{N}_{d}(0,\tilde{\Sigma}_{n})$. Generating a draw from $\mathsf{N}_{d}(0,\tilde{\Sigma}_{n})$ via this multiplier bootstrap may be numerically preferable to first drawing from $\mathsf{N}_{d}(0,\mathsf{I}_{d})$ and then premultiplying this by $\tilde{\Sigma}_{n}^{1/2}$ since the calculation of the matrix square root may be costly for $d$ large.
4 Normalized winsorized means
In practice one often normalizes the data by estimates of the $\sigma_{2,j}$ to bring the variables on the same scale. We now describe how the Gaussian and bootstrap approximations established so far remain valid upon normalization by the diagonal elements $\tilde{\sigma}_{n,j}=\tilde{\Sigma}_{n,j,j}^{1/2}$ of the robust estimator $\tilde{\Sigma}_{n}$ of $\Sigma$, cf. (8). Writing $D=\text{diag}(\sigma_{2,1},\ldots,\sigma_{2,d})$ and $\Sigma_{0}=D^{-1}\Sigma D^{-1}$ for the correlation matrix of the (centered) uncontaminated $X_{1},\ldots,X_{n}$, one has the following Gaussian approximation for the vector $S_{n,W,S}$ of normalized winsorized means with elements (we leave the quotients undefined if one of the variance estimators equals 0)
$S_{n,W,S,j}=\frac{1}{\sqrt{n}\tilde{\sigma}_{n,j}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}(\tilde{X}_{i,j})-\mu_{j}\bigr]},\qquad j=1,\ldots,d.$ (11)
Theorem 4.1.
Let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n},\varepsilon_{n}^{\prime}\in(0,1/2)$, with $\varepsilon_{n}$ as in (4) and $\varepsilon_{n}^{\prime}$ as in (7), then for $Z^{\prime}\sim\mathsf{N}_{d}(0,\Sigma_{0})$,
$\rho_{n,W,S}\mathrel{\mathop{\ordinarycolon}}$ $=\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(S_{n,W,S}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z^{\prime}\in H\bigr)}\Bigr\rvert}$
$\leq C\mathinner{\Biggl(\mathfrak{A}_{n}+\sqrt{\log(d)\log(dn)}\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\Biggr)},$ (12)
where $\mathfrak{A}_{n}$ is the upper bound on $\rho_{n,W}$ in (5) of Theorem 2.1 and $C$ is a constant depending only on $b_{1},b_{2},\lambda_{1},\lambda_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$ and $m$.
In particular, $\rho_{n,W,S}\to 0$ if $\sqrt{n\log(d)}\overline{\eta}_{n}^{1-\frac{1}{m}}\to 0$ and $\log(d)/n^{\frac{m-2}{5m-2}}\to 0$.
[…]
Theorem 4.2.
Let Assumption 2.1 be satisfied with $m>2$. Let $n>24$. If $\varepsilon_{n},\varepsilon_{n}^{\prime}\in(0,1/2)$, with $\varepsilon_{n}$ as in (4) and $\varepsilon_{n}^{\prime}$ as in (7), and $\tilde{Z}^{\prime}\sim\mathsf{N}_{d}(0,\tilde{\Sigma}_{n,0})$ conditionally on $\tilde{X}_{1},\ldots,\tilde{X}_{n}$, it holds with probability at least $1-\frac{24}{n}$ that
$\tilde{\rho}_{n,W,S}\mathrel{\mathop{\ordinarycolon}}$ $=\sup_{H\in\mathcal{H}}\mathinner{\!\biggl\lvert\mathbb{P}\mathinner{\bigl(S_{n,W,S}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(\tilde{Z}^{\prime}\in H\mid\tilde{X}_{1},\ldots,\tilde{X}_{n}\bigr)}\biggr\rvert}$
$\leq\mathfrak{B}_{n}+C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\Biggr)}^{1/2},$ (13)
where $\mathfrak{B}_{n}$ is the upper bound on $\rho_{n,W,S}$ in (4.1) of Theorem 4.1 and $C$ is a constant depending only on $b_{1},b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$ and $m$.
In particular, $\tilde{\rho}_{n,W,S}\to 0$ in probability if $\sqrt{n\log(d)}\overline{\eta}_{n}^{1-\frac{1}{m}}\to 0$ and $\log(d)/n^{\frac{m-2}{5m-2}}\to 0$.
5 Trimmed means
In this section we show that the Gaussian and bootstrap approximations established so far for winsorized means carry over to (suitably) trimmed means allowing for exactly the same growth rates of $\overline{\eta}_{n}$ and $d$. For $\varepsilon_{n}$ as in (4), let $I_{n}\mathrel{\mathop{\ordinarycolon}}=\mathinner{\bigl\{\lceil\varepsilon_{n}n\rceil,\ldots,\lceil(1-\varepsilon_{n})n\rceil\bigr\}}$ with cardinality
$|I_{n}|=\lceil(1-\varepsilon_{n})n\rceil-\lceil\varepsilon_{n}n\rceil+1=n-\lfloor\varepsilon_{n}n\rfloor-\lceil\varepsilon_{n}n\rceil+1,$
and consider the vector of trimmed means $S_{n,T}\in\mathbb{R}^{d}$ with entries ($j=1,\ldots,d$)
$S_{n,T,j}=\frac{\sqrt{n}}{|I_{n}|}\sum_{i\in I_{n}}\mathinner{\bigl[\tilde{X}_{i,j}^{*}-\mu_{j}\bigr]}=\frac{\sqrt{n}}{n-\lfloor\varepsilon_{n}n\rfloor-\lceil\varepsilon_{n}n\rceil+1}\sum_{i=\lceil\varepsilon_{n}n\rceil}^{\lceil(1-\varepsilon_{n})n\rceil}\mathinner{\bigl[\tilde{X}_{i,j}^{*}-\mu_{j}\bigr]}.$ (14)
Analogously to the winsorized means in (3), the trimmed mean in (14) can be implemented in a fully data-driven way: Its amount of trimming is governed by $\varepsilon_{n}$ in (4), which does not depend on any unknown population quantities. [^5] As is typical in the literature on inference under data contamination, an upper bound $\overline{\eta}_{n}$ on the contamination rate must be supplied, however. This sets this trimmed mean apart from the one studied in Theorem 2 in Resende (2024), where the amount of trimming depends on $m$. Furthermore, the bound in Theorem 5.1 below is valid over the larger family of sets $\mathcal{H}$ instead of $\mathcal{R}$. In analogy to $\rho_{n,W}$ define
$\rho_{n,T}=\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(S_{n,T}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z\in H\bigr)}\Bigr\rvert}$.
The following theorem is the trimmed mean counterpart to Theorem 2.1.
Theorem 5.1.
Let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n}\in(0,1/2)$, with $\varepsilon_{n}$ as in (4), then
$\rho_{n,T}\leq\mathfrak{A}_{n}+C\sqrt{n\log(d)}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr)},$ (15)
where $\mathfrak{A}_{n}$ is the upper bound on $\rho_{n,W}$ in (5) of Theorem 2.1 and $C$ is a constant depending only on $b_{1},b_{2},\lambda_{1},\lambda_{2}$ and $m$.
In particular, $\rho_{n,T}\to 0$ if $\sqrt{n\log(d)}\overline{\eta}_{n}^{1-\frac{1}{m}}\to 0$ and $\log(d)/n^{\frac{m-2}{5m-2}}\to 0$.
Theorem 5.1 is proven by establishing that $S_{n,T}$ is sufficiently close to $S_{n,W}$ in the supremum-norm, in order to use a Gaussian anti-concentration inequality to deduce (15) from (5) in Theorem 2.1. This explains the presence of the additional summand on the right-hand side in (15). The same discussion as that following Theorem 2.1 also applies to the trimmed mean in (14).
The following theorem is the trimmed mean analogue to Theorem 3.2 on bootstrap approximation to the distribution of winsorized means.
Theorem 5.2.
Let Assumption 2.1 be satisfied with $m>2$. Let $n>24$. If $\varepsilon_{n},\varepsilon_{n}^{\prime}\in(0,1/2)$, with $\varepsilon_{n}$ as in (4) and $\varepsilon_{n}^{\prime}$ as in (7), and $\tilde{Z}\sim\mathsf{N}_{d}(0,\tilde{\Sigma}_{n})$ conditionally on $\tilde{X}_{1},\ldots,\tilde{X}_{n}$, it holds with probability at least $1-\frac{24}{n}$ that
$\tilde{\rho}_{n,T}\mathrel{\mathop{\ordinarycolon}}$ $=\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(S_{n,T}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(\tilde{Z}\in H\mid\tilde{X}_{1},\ldots,\tilde{X}_{n}\bigr)}\Bigr\rvert}$
$\leq\mathfrak{C}_{n}+C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\Biggr)}^{1/2},$
where $\mathfrak{C}_{n}$ is the upper bound on $\rho_{n,T}$ in (15) of Theorem 5.1 and $C$ is a constant depending only on $b_{1},b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$ and $m$.
In particular, $\tilde{\rho}_{n,T}\to 0$ in probability if $\sqrt{n\log(d)}\overline{\eta}_{n}^{1-\frac{1}{m}}\to 0$ and $\log(d)/n^{\frac{m-2}{5m-2}}\to 0$.
Theorem 5.2 allows for exactly the same growth rates of $\overline{\eta}_{n}$ and $d$ as Theorem 3.2 for winsorized means. Note that we have chosen to keep the estimator $\tilde{\Sigma}_{n}$, which is based on winsorized means. Of course, one could also use a trimmed mean based estimator obeying the same performance guarantees (in fact, any estimator obeying the same performance guarantees as $\tilde{\Sigma}_{n}$ in (10) suffices, cf. also the discussion following (6)).
6 Numerical results
The P-P (probability-probability) plots in Figures 2 and 3 below illustrate the theoretical results of the paper. To allow a comparison to existing Gaussian approximations to the distribution of $S_{n}=n^{-1/2}\sum_{i=1}^{n}(X_{i}-\mu)$, which all are for uncontaminated $X_{i}$, we set $\overline{\eta}_{n}=0$ throughout. The $X_{i,j}$ are i.i.d., following a $t(\nu)$-distribution with $\nu$ degrees of freedom such that $\mu=0$ and $\Sigma=\text{diag}(\nu/(\nu-2),\ldots,\nu/(\nu-2))$. We consider $\nu=3.01$ (Figure 2) and $\nu=4.01$ (Figure 3), corresponding to the $X_{i,j}$ having all moments strictly less than $3.01$ and $4.01$, respectively. Throughout, $n=200$ and we study $d\in\mathinner{\{500,5{,}000\}}$. When implementing $S_{n,W}$ and $S_{n,T}$$\lambda_{1}$ and $\lambda_{1}^{\prime}$ are irrelevant as $\overline{\eta}_{n}=0$. We set $\lambda_{2}=0.1$ throughout. We implement $\tilde{\Sigma}_{n}$ with $\lambda_{2}^{\prime}=0.07$ since then the number of observations winsorized in implementing $\tilde{\Sigma}_{n}$ is typically the same as in implementing $S_{n,W}$ and $S_{n,T}$, i.e., $\varepsilon_{n}n=\lambda_{2}\log(dn)\approx\lambda_{2}^{\prime}\log(d^{2}n)=\varepsilon_{n}^{\prime}n$, which is found to work well. All results are based on 10,00010{,}000 replications and we write $x||_{\infty}=\max_{1\leq j\leq d}|x_{j}|$ for $x\in\mathbb{R}^{d}$.
[…]
Figure 3 increases $\nu$ to $4.01$ from $3.01$. As expected, this generally improves the performance of all approximations. In conclusion, the Gaussian and bootstrap approximations to the winsorized mean work well in all settings considered. We note, however, that whenever $n\varepsilon_{n}$ or $n\varepsilon_{n}^{\prime}$ cross an integer value, the number of winsorized observations changes, which can lead to a “discontinuity” in the performance. We also emphasize again that we have investigated scenarios without contamination. The normal and bootstrap approximations for the sample mean break down completely under contamination.
$\nu=3.01$$n=200$ and $d=500$
$\nu=3.01$$n=200$ and d=5,000d=5{,}000
[…]
$\nu=4.01$$n=200$ and $d=500$
$\nu=4.01$$n=200$ and d=5,000d=5{,}000
[…]
Appendix A A useful decomposition
We begin by outlining a decomposition of $S_{n,W}$ used in the proof of Theorem 2.1 along with some further notation used in the appendix. For $p\in(0,1)$ and a random variable $X$, denote by $Q_{p}(X)$ the $p$-quantile of the distribution of $X$, that is
$Q_{p}(X)=\inf\mathinner{\bigl\{z\in\mathbb{R}\mathrel{\mathop{\ordinarycolon}}\mathbb{P}(X\leq z)\geq p\bigr\}}.$
For $A_{+}=1-\lambda_{1}^{-1}\mathds{1}\mathinner{\{\overline{\eta}_{n}>0\}}$ and $A_{-}=1+\lambda_{1}^{-1}\mathds{1}\mathinner{\{\overline{\eta}_{n}>0\}}$, define (recall $\varepsilon_{n}$ from (4))
$c_{1,n}\mathrel{\mathop{\ordinarycolon}}=-A_{+}W_{0}\mathinner{\bigl(-e^{-(\frac{\log(dn)}{\varepsilon_{n}n}+A_{+})/A_{+}}\bigr)}\in(0,A_{+})$ (A.1)
as well as
$c_{2,n}\mathrel{\mathop{\ordinarycolon}}=-A_{-}W_{-1}\mathinner{\bigl(-e^{-(\frac{\log(dn)}{\varepsilon_{n}n}+A_{-})/A_{-}}\bigr)}\in(A_{-},\infty);$ (A.2)
here $W_{0}$ and $W_{-1}$ denote the principal and lower branch, respectively, of Lambert’s $W$ function (cf., e.g., Corless et al. (1996)). Because $A_{+}\leq A_{-}$, note that $0<c_{1,n}<c_{2,n}<\infty$. Next, for $j=1,\ldots,d$, let
$\underline{\alpha}_{j}\mathrel{\mathop{\ordinarycolon}}=Q_{c_{1,n}\varepsilon_{n}}(X_{1,j})\qquad\text{and}\qquad\overline{\alpha}_{j}\mathrel{\mathop{\ordinarycolon}}=Q_{c_{2,n}\varepsilon_{n}}(X_{1,j})$ (A.3)
as well as
$\underline{\beta}_{j}\mathrel{\mathop{\ordinarycolon}}=Q_{1-c_{2,n}\varepsilon_{n}}(X_{1,j})\qquad\text{and}\qquad\overline{\beta}_{j}\mathrel{\mathop{\ordinarycolon}}=Q_{1-c_{1,n}\varepsilon_{n}}(X_{1,j}),$ (A.4)
In the appendix only we shall sometimes impose
$\mathfrak{T}_{n}^{(1)}\mathrel{\mathop{\ordinarycolon}}=2\varepsilon_{n}+\frac{\log(dn)}{n}+\sqrt{\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{2}+4\frac{\log(dn)}{n}\varepsilon_{n}}<\frac{1}{2},$ (A.5)
which, however, is not imposed for any of our Gaussian and bootstrap approximations in this paper, as these will be argued to hold trivially when (A.5) is not satisfied. By Lemma G.2 one has that (A.5) ensures that
$0<\varepsilon_{n}\min(c_{1,n},c_{2,n})\leq\varepsilon_{n}(c_{1,n}+c_{2,n})<1/2,$
implying that $\underline{\alpha}_{j}\leq\underline{\beta}_{j}$ and $\overline{\alpha}_{j}\leq\overline{\beta}_{j}$, such that the decompositions in (A.7) and (A.8) below are well-defined.
By definition, $\hat{\alpha}_{j}=\tilde{X}_{\lceil\varepsilon_{n}n\rceil,j}^{*}$ and $\hat{\beta}_{j}=\tilde{X}_{\lceil(1-\varepsilon_{n})n\rceil,j}^{*}$. Lemma G.3 and the union bound, together with obvious monotonicity properties of $(a,b)\mapsto\phi_{a,b}$, show that with probability at least $1-\frac{4}{n}$ we have simultaneously for $j=1,\ldots,d$ (the inequalities $\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}\leq\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}\leq\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}$ and hence)
$\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(\tilde{X}_{i,j})-\mu_{j}\bigr]}\leq\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}(\tilde{X}_{i,j})-\mu_{j}\bigr]}\leq\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(\tilde{X}_{i,j})-\mu_{j}\bigr]}.$ (A.6)
The far right-hand of side of (A.6) can be decomposed as
$\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(\tilde{X}_{i,j})-\mu_{j}\bigr]}$ $=\underbrace{\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(\tilde{X}_{i,j})-\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})\bigr]}}_{\overline{I}_{n,j,1}}$
$+\underbrace{\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})\bigr]}}_{\overline{I}_{n,j,2}}$
$+\underbrace{\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})-\mu_{j}\bigr]}}_{\overline{I}_{n,j,3}}.$ (A.7)
Similarly, the left-hand side of (A.6) can be decomposed as
$\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(\tilde{X}_{i,j})-\mu_{j}\bigr]}=\underline{I}_{n,j,1}+\underline{I}_{n,j,2}+\underline{I}_{n,j,3},$ (A.8)
with $\underline{I}_{n,j,k}$, $j=1,\ldots,d,\ k=1,2,3$, defined analogously to the $\overline{I}_{n,j,k}$. Define
$\overline{I}_{n,k}=\max_{j=1,\ldots,d}\mathinner{\!\bigl\lvert\overline{I}_{n,j,k}\bigr\rvert}\qquad\text{and}\qquad\underline{I}_{n,k}=\max_{j=1,\ldots,d}\mathinner{\!\bigl\lvert\underline{I}_{n,j,k}\bigr\rvert},\qquad k=1,3,$ (A.9)
as well as
$\overline{I}_{n,2}=\mathinner{\bigl(\overline{I}_{n,1,2},\ldots,\overline{I}_{n,d,2}\bigr)}^{\prime}\qquad\text{and}\qquad\underline{I}_{n,2}=\mathinner{\bigl(\underline{I}_{n,1,2},\ldots,\underline{I}_{n,d,2}\bigr)}^{\prime}.$
Throughout, $\overline{\mathbb{R}}=\mathbb{R}\cup\mathinner$$\{$-$\infty$$\infty\}$}. Thus, writing $Y_{n,j}=n^{-1/2}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}(\tilde{X}_{i,j})-\mu_{j}\bigr]}$ and $Y_{n}=\mathinner{\bigl(Y_{n,1},\ldots,Y_{n,d}\bigr)}^{\prime}$, one has with probability at least $1-\frac{4}{n}$ that
$\sqrt{n}(\underline{I}_{n,2}-\underline{I}_{n,1}-\underline{I}_{n,3})\leq Y_{n}\leq\sum_{k=1}^{3}\sqrt{n}\overline{I}_{n,k},$ (A.10)
where here and in the sequel (i) all inequalities between vectors are understood elementwise, and (ii) with some abuse of notation we define, for a vector $x\in\overline{\mathbb{R}}^{m}$, say, and a real number $a$, the sum $x+a$ coordinate-wise as $(x_{1}+a,\ldots,x_{m}+a)^{\prime}$ , with the usual convention that $\infty+a=\infty$ and $\infty+a=-\infty$. In the next section we study the left- and right-hand sides of the previous display.
Appendix B Preparatory lemmas
Let $\varepsilon_{n}$ be as in (4) and recall the quantities defined in (A.3) and (A.4). Let $\overline{\Sigma}_{\varepsilon_{n}}$ and $\underline{\Sigma}_{\varepsilon_{n}}$ be the matrices with $(j,k)$th entry
$\overline{\Sigma}_{\varepsilon_{n},j,k}=\mathbb{E}\mathinner{\Bigl[\mathinner{\bigl(\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})\bigr)}\mathinner{\bigl(\phi_{\overline{\alpha}_{k},\overline{\beta}_{k}}(X_{1,k})-\mathbb{E}\phi_{\overline{\alpha}_{k},\overline{\beta}_{k}}(X_{1,k}\bigr)}\Bigr]}$
and
$\underline{\Sigma}_{\varepsilon_{n},j,k}=\mathbb{E}\mathinner{\Bigl[\mathinner{\bigl(\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{1,j})-\mathbb{E}\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{1,j})\bigr)}\mathinner{\bigl(\phi_{\underline{\alpha}_{k},\underline{\beta}_{k}}(X_{1,k})-\mathbb{E}\phi_{\underline{\alpha}_{k},\underline{\beta}_{k}}(X_{1,k}\bigr)}\Bigr]},$
respectively. The following lemma bounds the distance of these covariance matrices to the covariance matrix $\Sigma$ of the vector $X_{1}$.
Lemma B.1.
Let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n}$ is chosen as in (4) satisfying (A.5), then, for a positive constant $C$ depending only on $\lambda_{2}$ and $m$, it holds that
$\max_{j,k\in\mathinner{\{1,\ldots,d\}}}\mathinner{\Bigl(\mathinner{\!\bigl\lvert\overline{\Sigma}_{\varepsilon_{n},j,k}-\Sigma_{j,k}\bigr\rvert}\vee\mathinner{\!\bigl\lvert\underline{\Sigma}_{\varepsilon_{n},j,k}-\Sigma_{j,k}\bigr\rvert}\Bigr)}\leq C\sigma_{m}^{2}\varepsilon_{n}^{1-\frac{2}{m}}=\mathrel{\mathop{\ordinarycolon}}\mathfrak{T}_{n}^{(2)}.$ (B.1)
Proof.
We only establish (B.1) for $\max_{j,k\in\mathinner{\{1,\ldots,d\}}}\mathinner{\!\bigl\lvert\overline{\Sigma}_{\varepsilon_{n},j,k}-\Sigma_{j,k}\bigr\rvert}$ as the proof is identical for $\max_{j,k\in\mathinner{\{1,\ldots,d\}}}\mathinner{\!\bigl\lvert\underline{\Sigma}_{\varepsilon_{n},j,k}-\Sigma_{j,k}\bigr\rvert}$.
Fix $1\leq j\leq d$ and note that
$\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})$
$=X_{1,j}+(\overline{\alpha}_{j}-X_{1,j})\mathds{1}\mathinner{\bigl(X_{1,j}<\overline{\alpha}_{j}\bigr)}+(\overline{\beta}_{j}-X_{1,j})\mathds{1}\mathinner{\bigl(X_{1,j}>\overline{\beta}_{j}\bigr)}$
$-\mathbb{E}\mathinner{\bigl(X_{1,j}+(\overline{\alpha}_{j}-X_{1,j})\mathds{1}\mathinner{\bigl(X_{1,j}<\overline{\alpha}_{j}\bigr)}+(\overline{\beta}_{j}-X_{1,j})\mathds{1}\mathinner{\bigl(X_{1,j}>\overline{\beta}_{j}\bigr)}\bigr)}$
$=\mathinner{\bigl[X_{1,j}-\mathbb{E}X_{1,j}\bigr]}+\mathinner{\bigl[(\overline{\alpha}_{j}-X_{1,j})\mathds{1}\mathinner{\bigl(X_{1,j}<\overline{\alpha}_{j}\bigr)}-\mathbb{E}(\overline{\alpha}_{j}-X_{1,j})\mathds{1}\mathinner{\bigl(X_{1,j}<\overline{\alpha}_{j}\bigr)}\bigr]}$
$+\mathinner{\bigl[(\overline{\beta}_{j}-X_{1,j})\mathds{1}\mathinner{\bigl(X_{1,j}>\overline{\beta}_{j}\bigr)}-\mathbb{E}(\overline{\beta}_{j}-X_{1,j})\mathds{1}\mathinner{\bigl(X_{1,j}>\overline{\beta}_{j}\bigr)}\bigr]}$
$=\mathrel{\mathop{\ordinarycolon}}T_{1,j}+T_{2,j}+T_{3,j}.$
Thus, for $j,k\in\mathinner{\{1,\ldots,d\}}$, it follows from the Cauchy-Schwarz inequality that
$\mathinner{\!\bigl\lvert\overline{\Sigma}_{\varepsilon_{n},j,k}-\Sigma_{j,k}\bigr\rvert}$ $=\mathinner{\!\bigl\lvert\mathbb{E}(T_{1,j}+T_{2,j}+T_{3,j})(T_{1,k}+T_{2,k}+T_{3,k})-\mathbb{E}(T_{1,j}T_{1,k})\bigr\rvert}$
$\leq|\mathbb{E}T_{1,j}T_{2,k}|+|\mathbb{E}T_{1,j}T_{3,k}|+|\mathbb{E}T_{2,j}T_{1,k}|+|\mathbb{E}T_{2,j}T_{2,k}|$
$+|\mathbb{E}T_{2,j}T_{3,k}|+|\mathbb{E}T_{3,j}T_{1,k}|+|\mathbb{E}T_{3,j}T_{2,k}|+|\mathbb{E}T_{3,j}T_{3,k}|$
$\leq|\mathbb{E}T_{1,j}T_{2,k}|+|\mathbb{E}T_{1,j}T_{3,k}|+|\mathbb{E}T_{2,j}T_{1,k}|+|\mathbb{E}T_{3,j}T_{1,k}|$
$+\mathinner{\bigl(\mathbb{E}T_{2,j}^{2}\mathbb{E}T_{2,k}^{2}\bigr)}^{1/2}+\mathinner{\bigl(\mathbb{E}T_{2,j}^{2}\mathbb{E}T_{3,k}^{2}\bigr)}^{1/2}+\mathinner{\bigl(\mathbb{E}T_{3,j}^{2}\mathbb{E}T_{2,k}^{2}|\bigr)}^{1/2}+\mathinner{\bigl(\mathbb{E}T_{3,j}^{2}\mathbb{E}T_{3,k}^{2}\bigr)}^{1/2}.$ (B.2)
We proceed by bounding the far right-hand side of (B.2). To this end, note first that for all $j=1,\ldots,d$ it holds that
$\mathbb{E}T_{2,j}^{2}$ $=\mathbb{E}\mathinner{\bigl[(\overline{\alpha}_{j}-X_{1,j})\mathds{1}\mathinner{\bigl(X_{1,j}<\overline{\alpha}_{j}\bigr)}-\mathbb{E}(\overline{\alpha}_{j}-X_{1,j})\mathds{1}\mathinner{\bigl(X_{1,j}<\overline{\alpha}_{j}\bigr)}\bigr]}^{2}$
$\leq\mathbb{E}(\overline{\alpha}_{j}-X_{1,j})^{2}\mathds{1}\mathinner{\bigl(X_{1,j}<\overline{\alpha}_{j}\bigr)}.$
Recall that $\overline{\alpha}_{j}=Q_{c_{2,n}\varepsilon_{n}}(X_{1,j})$ such that $\mathbb{P}(X_{1,j}<\overline{\alpha}_{j})\leq c_{2,n}\varepsilon_{n}$. Thus, it follows by Lemma G.1, followed by Hölder’s inequality, that there exists a positive constant $C$ depending only on $\lambda_{2}$ and $m$ such that
$\mathbb{E}T_{2,j}^{2}$ $\leq\mathbb{E}(\mu_{j}+\sigma_{m}/(1-c_{2,n}\varepsilon_{n})^{1/m}-X_{1,j})^{2}\mathds{1}\mathinner{\bigl(X_{1,j}<\overline{\alpha}_{j}\bigr)}$
$\leq\mathinner{\Bigl(\mathbb{E}\mathinner{\!\bigl\lvert\mu_{j}-X_{1,j}+\sigma_{m}/(1-c_{2,n}\varepsilon_{n})^{1/m}\bigr\rvert}^{m}\Bigr)}^{2/m}\mathbb{P}(X_{1,j}<\overline{\alpha}_{j})^{1-\frac{2}{m}}$
$\leq\mathinner{\Bigl(2^{m}\sigma_{m}^{m}+2^{m}\sigma_{m}^{m}/(1-c_{2,n}\varepsilon_{n})\Bigr)}^{2/m}(c_{2,n}\varepsilon_{n})^{1-\frac{2}{m}}$
$\leq C\sigma_{m}^{2}\varepsilon_{n}^{1-\frac{2}{m}},$ (B.3)
the last inequality using that $c_{2,n}\varepsilon_{n}<1/2$ by Lemma G.2 and (A.5) and the same lemma also bounding $c_{2,n}$ from above in terms of $\lambda_{2}$ only.
By the same arguments, using $\mathbb{P}(X_{1,j}>\overline{\beta}_{j})=1-\mathbb{P}(X_{1,j}\leq\overline{\beta}_{j})\leq c_{1,n}\varepsilon_{n}$ and adjusting $C$, one also has that
$\mathbb{E}T_{3,j}^{2}\leq C\sigma_{m}^{2}\varepsilon_{n}^{1-\frac{2}{m}}.$ (B.4)
Furthermore,
$\mathbb{E}T_{1,j}T_{2,k}$ $=\mathbb{E}\mathinner{\Bigl([X_{1,j}-\mu_{j}]\mathinner{\bigl[(\overline{\alpha}_{k}-X_{1,k})\mathds{1}\mathinner{\bigl(X_{1,k}<\overline{\alpha}_{k}\bigr)}-\mathbb{E}(\overline{\alpha}_{k}-X_{1,k})\mathds{1}\mathinner{\bigl(X_{1,k}<\overline{\alpha}_{k}\bigr)}\bigr]}\Bigr)}$
$=\mathbb{E}\mathinner{\bigl([X_{1,j}-\mu_{j}](\overline{\alpha}_{k}-X_{1,k})\mathds{1}\mathinner{\bigl(X_{1,k}<\overline{\alpha}_{k}\bigr)}\bigr)}.$
Recallling that $\overline{\alpha}_{j}=Q_{c_{2,n}\varepsilon_{n}}(X_{1,j})$, it follows by Lemma G.1 and Hölder’s inequality, that
$\mathinner{\!\bigl\lvert\mathbb{E}T_{1,j}T_{2,k}\bigr\rvert}$
$\leq\mathbb{E}\mathinner{\bigl(|X_{1,j}-\mu_{j}||\mu_{k}+\sigma_{m}/(1-c_{2,n}\varepsilon_{n})^{1/m}-X_{1,k}|\mathds{1}\mathinner{\bigl(X_{1,k}<\overline{\alpha}_{k}\bigr)}\bigr)}$
$\leq\mathbb{E}\mathinner{\bigl(|X_{1,j}-\mu_{j}||X_{1,k}-\mu_{k}|+|X_{1,j}-\mu_{j}|\sigma_{m}/(1-c_{2,n}\varepsilon_{n})^{1/m}\bigr)}\mathds{1}\mathinner{\bigl(X_{1,k}<\overline{\alpha}_{k}\bigr)}$
$\leq\mathinner{\Bigl(\mathbb{E}\mathinner{\bigl(|X_{1,j}-\mu_{j}||X_{1,k}-\mu_{k}|+|X_{1,j}-\mu_{j}|\sigma_{m}/(1-c_{2,n}\varepsilon_{n})^{1/m}\bigr)}^{\frac{m}{2}}\Bigr)}^{\frac{2}{m}}(c_{2,n}\varepsilon_{n})^{1-\frac{2}{m}}$
$\leq\mathinner{\Bigl(2^{m/2}(\sigma_{m}^{m}+\sigma_{m/2}^{m/2}\sigma_{m}^{m/2}/(1-c_{2,n}\varepsilon_{n})^{1/2}\Bigr)}^{2/m}(c_{2,n}\varepsilon_{n})^{1-\frac{2}{m}}$
$\leq C\sigma_{m}^{2}\varepsilon_{n}^{1-\frac{2}{m}},$ (B.5)
the last inequality using that $c_{2,n}\varepsilon_{n}<1/2$ by Lemma G.2 and (A.5) and the same lemma also bounding $c_{2,n}$ from above in terms of $\lambda_{2}$ only (adjust $C$ if needed).
By the same arguments,
$\mathinner{\!\bigl\lvert\mathbb{E}T_{1,j}T_{3,k}\bigr\rvert}\leq C\sigma_{m}^{2}\varepsilon_{n}^{1-\frac{2}{m}}.$ (B.6)
Therefore, observing that none of the upper bounds in (B.3)–(B.6) depend on $j$ and $k$ it follows that (adjust $C$ if needed)
$\max_{j,k\in\mathinner{\{1,\ldots,d\}}}\mathinner{\!\bigl\lvert\overline{\Sigma}_{\varepsilon_{n},j,k}-\Sigma_{j,k}\bigr\rvert}\leq C\sigma_{m}^{2}\varepsilon_{n}^{1-\frac{2}{m}}.$
The following lemma applies a high-dimensional Gaussian approximation to $\underline{I}_{n,2}$ and $\overline{I}_{n,2}$, which are both sums of winsorized, and hence bounded, random variables.
Lemma B.2.
Let Assumption 2.1 be satisfied with $m>2$. Let $\varepsilon_{n}$ be chosen as in (4) satisfying (A.5) and suppose that there exists a strictly positive constants $\mathfrak{b}_{1}$ such that for all $j=1,\ldots,d$
$\mathbb{E}\mathinner{\bigl[\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{1,j})-\mathbb{E}\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{1,j})\bigr]}^{2}>\mathfrak{b}_{1}^{2}\qquad\text{and}\qquad\mathbb{E}\mathinner{\bigl[\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})\bigr]}^{2}>\mathfrak{b}_{1}^{2}.$
Then, for $\overline{Z}\sim\mathsf{N}_{d}(0,\overline{\Sigma}_{\varepsilon_{n}})$ and $\underline{Z}\sim\mathsf{N}_{d}(0,\underline{\Sigma}_{\varepsilon_{n}})$,
$\sup_{H\in\mathcal{H}}\mathinner{\!\biggl\lvert\mathbb{P}\mathinner{\biggl(\frac{1}{n^{1/2}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})\bigr]}\in H\biggr)}-\mathbb{P}\mathinner{\bigl(\overline{Z}\in H\bigr)}\biggr\rvert}$
and
$\sup_{H\in\mathcal{H}}\mathinner{\!\biggl\lvert\mathbb{P}\mathinner{\biggl(\frac{1}{n^{1/2}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{i,j})-\mathbb{E}\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{i,j})\bigr]}\in H\biggr)}-\mathbb{P}\mathinner{\bigl(\underline{Z}\in H\bigr)}\biggr\rvert}$
are bounded from above by
$C\sigma_{m}^{1/2}\mathinner{\biggl(\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr)}^{\frac{1}{4}},$
where $C$ is a constant depending only on $\mathfrak{b}_{1}$$b_{2}$$\lambda_{1}$$\lambda_{2}$, and $m$.
Proof.
We establish the first bound by verifying the conditions of Theorem 2.1 in Chernozhukov et al. (2022), cf. in particular the display following that theorem. The second bound is proven analogously. First, note that for $j=1,\ldots,d$
$\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})\leq\overline{\beta}_{j}-\overline{\alpha}_{j}=Q_{1-c_{1,n}\varepsilon_{n}}(X_{1,j})-Q_{c_{2,n}\varepsilon_{n}}(X_{1,j}),$
Furthermore, by Lemma G.1 and $\varepsilon_{n}\geq\lambda_{2}\frac{\log(dn)}{n}$, the far right-hand side of the previous display is upper bounded by
$\frac{\sigma_{m}}{(c_{1,n}\varepsilon_{n})^{1/m}}+\frac{\sigma_{m}}{(c_{2,n}\varepsilon_{n})^{1/m}}\leq C_{1}\sigma_{m}\mathinner{\biggl(\frac{n}{\log(dn)}\biggr)}^{1/m},$
where, by Lemma G.2, $C_{1}$ is a constant depending only on $\lambda_{1}$$\lambda_{2}$ and $m$. Similarly,
$\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})$ $\geq\overline{\alpha}_{j}-\overline{\beta}_{j}=Q_{c_{2,n}\varepsilon_{n}}(X_{1,j})-Q_{1-c_{1,n}\varepsilon_{n}}(X_{1,j})$
$\geq-C_{1}\sigma_{m}\mathinner{\biggl(\frac{n}{\log(dn)}\biggr)}^{1/m},$
and we define
$B_{n}=2C_{1}\sigma_{m}\mathinner{\biggl(\frac{n}{\log(dn)}\biggr)}^{1/m}.$
By construction, $\mathinner{\!\bigl\lvert\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})\bigr\rvert}\leq\frac{1}{2}B_{n}$ for all $i=1,\ldots,n$ and $j=1,\ldots,d$. Furthermore, by assumption, $\mathbb{E}\mathinner{\bigl[\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})\bigr]}^{2}>\mathfrak{b}_{1}^{2}$, as well as
$\mathbb{E}\mathinner{\bigl[\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})\bigr]}^{4}$ $\leq B_{n}^{2}\mathbb{E}\mathinner{\bigl[\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})\bigr]}^{2}$
$\leq B_{n}^{2}\mathbb{E}\mathinner{\bigl[X_{1,j}-\mathbb{E}(X_{1,j})\bigr]}^{2}\leq B_{n}^{2}\sigma_{2}^{2}\leq B_{n}^{2}b_{2}^{2};$
a proof of the second inequality can be found in, e.g., Corollary 3 in Chow and Studden (1969). Therefore, by Theorem 2.1 and the display following it in Chernozhukov et al. (2022), there exists a constant $C_{2}$ depending only on $\mathfrak{b}_{1}$ and $b_{2}$ such that
$\sup_{H\in\mathcal{H}}\mathinner{\!\biggl\lvert\mathbb{P}\mathinner{\biggl(\frac{1}{n^{1/2}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})\bigr]}\in H\biggr)}-\mathbb{P}\mathinner{\bigl(\overline{Z}\in H\bigr)}\biggr\rvert}$
$\leq$ $C_{2}\mathinner{\biggl(\frac{B_{n}^{2}\log^{5}(dn)}{n}\biggr)}^{\frac{1}{4}}=C_{2}\cdot(2C_{1})^{1/2}\sigma_{m}^{1/2}\mathinner{\biggl(\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr)}^{\frac{1}{4}},$
from which the conclusion of the lemma follows. ∎
Lemma B.3.
Let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n}$ is chosen as in (4) and satisfies (A.5), then for $\overline{I}_{n,3}$ and $\underline{I}_{n,3}$ as defined in (A.9), it holds that
$\underline{I}_{n,3}\vee\overline{I}_{n,3}\leq\sigma_{m}C\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr)},$
for a positive constant $C$ depending only on $\lambda_{1}$$\lambda_{2}$, and $m$.
Proof.
This is a trivial consequence of Lemma G.5 and subadditivity of $z\mapsto z^{1-\frac{1}{m}}$. ∎
The following lemma collects some simple limits for later reference.
Lemma B.4.
Let $m>2$ and assume that $\sqrt{n\log(d)}\overline{\eta}_{n}^{1-\frac{1}{m}}\to 0$ as well as $\log(d)/n^{\frac{m-2}{5m-2}}\to 0$. Then
$\log^{2}(d)\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{2}{m}}\biggr)}\to 0,$ (B.7)
$\mathinner{\biggl[\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr]}^{\frac{1}{4}}+\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr)}\sqrt{n\log(d)}\to 0,$ (B.8)
$\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\to 0,$ (B.9)
and
$\sqrt{\log(d)\log(dn)}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{(m/2)\wedge 2}}\biggr)}\to 0.$ (B.10)
Proof.
We begin by proving (B.7), to which end we first establish that $\log^{2}(d)\overline{\eta}_{n}^{1-\frac{2}{m}}\to 0$. Since $\log(d)/n^{\frac{m-2}{5m-2}}\to 0$, one has that $\log^{2}(d)\overline{\eta}_{n}^{1-\frac{2}{m}}=o\mathinner{\bigl(n^{\frac{2m-4}{5m-2}}\overline{\eta}_{n}^{1-\frac{2}{m}}\bigr)}$ and
$n^{\frac{2m-4}{5m-2}}\overline{\eta}_{n}^{1-\frac{2}{m}}\to 0\iff n^{\frac{2m^{2}-4m}{(5m-2)(m-2)}}\overline{\eta}_{n}\to 0.$
The latter convergence is satisfied since $\sqrt{n}\overline{\eta}^{1-\frac{1}{m}}\to 0$ by assumption, which is equivalent to $n^{\frac{m}{2(m-1)}}\overline{\eta}_{n}\to 0$, and $n^{\frac{2m^{2}-4m}{(5m-2)(m-2)}}\overline{\eta}_{n}\leq n^{\frac{m}{2(m-1)}}\overline{\eta}_{n}\to 0$ for $m>2$.
Next,
$\log^{2}(d)\mathinner{\Bigl[\frac{\log(d)}{n}\Bigr]}^{1-\frac{2}{m}}=\frac{\log(d)^{3-\frac{2}{m}}}{n^{1-\frac{2}{m}}}\to 0\iff\frac{\log(d)}{n^{\frac{m-2}{3m-2}}}\to 0,$
the latter convergence following from $\log(d)/n^{\frac{m-2}{5m-2}}\to 0$ by assumption.
Finally,
$\log^{2}(d)\mathinner{\Bigl[\frac{\log(n)}{n}\Bigr]}^{1-\frac{2}{m}}\to 0$
by a standard subsequence argument, considering separately the cases of subsequences along which i) $n\leq d$ for which the convergence follows from the penultimate display and ii) $d<n$ for which the convergence follows from $m>2$. This establishes (B.7).
To prove (B.8), note that
$\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\to 0\iff\frac{\log(d)}{n^{\frac{m-2}{5m-2}}}\to 0,$
the latter convergence being true by assumption. Furthermore, $\sqrt{n\log(d)}\overline{\eta}_{n}^{1-\frac{1}{m}}\to 0$ by assumption and it remains to prove that
$\sqrt{n\log(d)}\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{m}}\to 0.$
To this end, note that
$\sqrt{n\log(d)}\mathinner{\Bigl(\frac{\log(d)}{n}\Bigr)}^{1-\frac{1}{m}}=\frac{\mathinner{\bigl[\log(d)\bigr]}^{\frac{3}{2}-\frac{1}{m}}}{n^{\frac{1}{2}-\frac{1}{m}}}\to 0\iff\frac{\log(d)}{n^{\frac{m-2}{3m-2}}}\to 0,$
where the latter convergence was already verified above. The convergence in the penultimate display now follows by considering separately subsequences along which $n\leq d$ and $d<n$, respectively (as in the concluding argument of the proof of (B.7)).
Next to establish, (B.9) note that we have already shown that $\log^{2}(d)\overline{\eta}_{n}^{1-\frac{2}{m}}\to 0$. Furthermore, in case $m\leq 4$,
$\log^{2}(d)\mathinner{\Bigl(\frac{\log(d)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}=\log^{2}(d)\mathinner{\Bigl(\frac{\log(d)}{n}\Bigr)}^{1-\frac{2}{m}}=\frac{\mathinner{\bigl[\log(d)\bigr]}^{3-\frac{2}{m}}}{n^{1-\frac{2}{m}}}\to 0\iff\frac{\log(d)}{n^{\frac{m-2}{3m-2}}}\to 0,$
the latter convergence being true because it is assumed that even $\log(d)/n^{\frac{m-2}{5m-2}}\to 0$. Conclude, once more, by a subsequence argument. The case of $m>4$ is handled analogously.
Finally, we establish (B.10) by showing that any subsequence possesses a further subsequence along which (B.10) is true. Thus, fix a subsequence. In case there exists a subsequence thereof along which $n\leq d$, it suffices to show that along this subsequence
$\log(d)\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(d)}{n}\Bigr]}^{1-\frac{1}{(m/2)\wedge 2}}\biggr)}\to 0,$
which is implied by (B.9).
In the remaining case (where there does not exists a further subsequence along which $n\leq d$), we have $d<n$ for $n$ sufficiently large, so that it suffices to show that
$\log(n)\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(n)}{n}\Bigr]}^{1-\frac{1}{(m/2)\wedge 2}}\biggr)}\to 0.$
Since $m>2$ this, in turn, is true if
$\log(n)\overline{\eta}_{n}^{1-\frac{2}{m}}=\sqrt{n}\overline{\eta}_{n}^{1-\frac{1}{m}}\cdot\frac{\log(n)}{\sqrt{n}}\overline{\eta}_{n}^{-\frac{1}{m}}\to 0,$
which, since $\sqrt{n}\overline{\eta}_{n}^{1-\frac{1}{m}}\to 0$ by assumption, is true in case $\frac{\log(n)}{\sqrt{n}}\overline{\eta}_{n}^{-\frac{1}{m}}$ is bounded. In case $\frac{\log(n)}{\sqrt{n}}\overline{\eta}_{n}^{-\frac{1}{m}}$ is unbounded, $\frac{\log(n)}{\sqrt{n}}\overline{\eta}_{n}^{-\frac{1}{m}}\geq 1$ along a further subsequence. The latter implies $\overline{\eta}_{n}\leq\frac{\log^{m}(n)}{n^{\frac{m}{2}}}$, such that
$\log(n)\overline{\eta}_{n}^{1-\frac{2}{m}}\leq\frac{[\log(n)]^{m-1}}{n^{\frac{m-2}{2}}}\to 0,$
because $m>2$. ∎
The following lemma shows that to establish Theorem 2.1, it suffices to prove it for the class of one-sided rectangles $H=\bigtimes_{j=1}^{d}[-\infty,t_{j}]$ with $t_{j}\in\overline{\mathbb{R}}=\mathbb{R}\cup\mathinner{\{-\infty,\infty\}}$ for $j=1,\ldots,d$.
Lemma B.5.
Suppose Theorem 2.1 holds for $\rho_{n,W}$ throughout replaced by
$\rho_{n,W}^{os}\mathrel{\mathop{\ordinarycolon}}=\sup_{t\in\overline{\mathbb{R}}^{d}}\mathinner{\!\bigl\lvert\mathbb{P}\mathinner{\bigl(S_{n,W}\leq t\bigr)}-\mathbb{P}\mathinner{\bigl(Z\leq t\bigr)}\bigr\rvert}$. (B.11)
Then, Theorem 2.1 holds (with a different constant $C$).
Proof.
Fix $a$ and $b$ in $\overline{\mathbb{R}}^{d}$ such that $\infty\leq a_{j}\leq b_{j}\leq\infty$ and let $H=\bigtimes_{j=1}^{d}[a_{j},b_{j}]\in\mathcal{H}$. Clearly,
$\mathbb{P}\mathinner{\bigl(S_{n,W}\in H\bigr)}=\mathbb{P}\mathinner{\bigl((S_{n,W}^{\prime},-S_{n,W}^{\prime})^{\prime}\leq(b^{\prime},-a^{\prime})^{\prime}\bigr)}.$ (B.12)
Furthermore, by (3)
$S_{n,W,j}=n^{-1/2}\sum_{i=1}^{n}\mathinner{\bigl(\phi_{-\hat{\beta}_{j},-\hat{\alpha}_{j}}(-\tilde{X}_{i,j})-(-\mu_{j})\bigr)},\qquad j=1,\ldots,d.$
Thus, the vectors $(X_{i}^{\prime},-X_{i}^{\prime})^{\prime}$ and $(\tilde{X}_{i}^{\prime},-\tilde{X}_{i}^{\prime})^{\prime}$ in $\mathbb{R}^{2d}$ satisfy Assumption 2.1 with $d$ there replaced by $2d$ (but with the same $m$, $b_{1}$$b_{2}$ and $\overline{\eta}_{n}$). Note that the covariance matrix of $(X_{1}^{\prime},-X_{1}^{\prime})^{\prime}$ is
$\Xi=\begin{pmatrix}\Sigma&-\Sigma\\ -\Sigma&\Sigma\end{pmatrix}$,
and let $Z_{2}\sim~\mathsf{N}_{2d}(0,\Xi)=\mathrel{\mathop{\ordinarycolon}}\nu^{\prime}$. Hence, the version of Theorem 2.1 that replaces $\rho_{n,W}$ by $\rho_{n,W}^{os}$ (and which is assumed to hold true) applies to $(S_{n,W}^{\prime},-S_{n,W}^{\prime})^{\prime}$ and yields, for $C^{*}$ a constant depending only on $b_{1},b_{2},\lambda_{1},\lambda_{2}$, and $m$, the upper bound
$C^{*}\mathinner{\Biggl(\mathinner{\biggl[\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr]}^{\frac{1}{4}}+\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr]}\sqrt{n\log(d)}\Biggr)}$
$+C^{*}\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{2}{m}}\biggr]}\Biggr)}^{1/2}$
on
$\mathinner{\!\bigl\lvert\mathbb{P}\mathinner{\bigl((S_{n,W},-S_{n,W})\leq(b^{\prime},-a^{\prime})^{\prime}\bigr)}-\mathbb{P}\mathinner{\bigl(Z_{2}\leq(b^{\prime},-a^{\prime})^{\prime}\bigr)}\bigr\rvert}$, (B.13)
Furthermore, for $\nu=\mathsf{N}_{d}(0,\Sigma)$$\nu^{\prime}$ is the image measure of $\nu$ under the mapping $\mathbb{R}^{d}\ni z\mapsto(z,-z)\in\mathbb{R}^{2d}$, that is $\nu^{\prime}=\nu\circ\mathinner{\bigl(z\mapsto(z,-z)\bigr)}^{-1}$. Thus,
$\mathbb{P}\mathinner{\bigl(Z_{2}\leq(b^{\prime},-a^{\prime})^{\prime}\bigr)}$ $=\nu^{\prime}\mathinner{\bigl([-\infty,b]\times[-\infty,-a]\bigr)}$
$=\nu\mathinner{\bigl(z\in\mathbb{R}^{d}\mathrel{\mathop{\ordinarycolon}}(z,-z)\in[-\infty,b]\times[-\infty,-a]\bigr)}$
$=\mathbb{P}\mathinner{\bigl((Z,-Z)\leq(b,-a)\bigr)}$
$=\mathbb{P}\mathinner{\bigl(Z\in H\bigr)}.$ (B.14)
Combining (B.12) and (B) with the upper bound in (B.13) obtained above (which does not depend on $a$ or $b$) delivers the claim. ∎
Appendix C Proof of Theorem 2.1
By Lemma B.5 it suffices to prove (5) for hyperractangles $H$ of the form $H=\bigtimes_{j=1}^{d}[-\infty,t_{j}]$ with $t_{j}\in\overline{\mathbb{R}}$. Furthermore, recalling the definitions of $\mathfrak{T}_{n}^{(1)}$ and $\mathfrak{T}_{n}^{2}$ in (A.5) and (B.1), respectively, and using that $\log(dn)/n\leq\lambda_{2}^{-1}\varepsilon_{n}$, one has with $C$ being the constant from (B.1) that
$\mathfrak{T}_{n}^{(1)}+\mathfrak{T}_{n}^{(2)}\leq 2\varepsilon_{n}+\lambda_{2}^{-1}\varepsilon_{n}+\sqrt{\lambda_{2}^{-2}\varepsilon_{n}^{2}+4\lambda_{2}^{-1}\varepsilon_{n}^{2}}+Cb_{2}^{2}\varepsilon_{n}^{1-\frac{2}{m}}\leq C_{1}\varepsilon_{n}^{1-\frac{2}{m}},$ (C.1)
for $C_{1}$ a constant depending only on $b_{2}$$\lambda_{2}$ and $m$, and where we used that $\varepsilon_{n}\leq\varepsilon_{n}^{1-\frac{2}{m}}$ as $\varepsilon_{n}\in(0,0.5)$. In the following we first suppose that
$C_{1}\varepsilon_{n}^{1-\frac{2}{m}}<\frac{(b_{1}^{2}\wedge 1)}{2},$ (C.2)
which, by the penultimate display, implies that $\mathfrak{T}_{n}^{(1)}<1/2$ such that (A.5) is satisfied. By Lemma G.4, there exists a constant $C_{2}$ depending only on $\lambda_{1}$$\lambda_{2}$, and $m$ such that
$\underline{I}_{n,1}\vee\overline{I}_{n,1}\leq C_{2}\sigma_{m}\overline{\eta}_{n}^{1-\frac{1}{m}}\mathrel{\mathop{\ordinarycolon}}=I_{n,1}$ (C.3)
with $I_{n,1}$ being non-random. Thus, from (A.10) it holds with probability at least $1-\frac{4}{n}$ that
$\underline{Y}_{n}\mathrel{\mathop{\ordinarycolon}}=\sqrt{n}\underline{I}_{n,2}-\sqrt{n}(I_{n,1}+\underline{I}_{n,3})\leq Y_{n}\leq\sqrt{n}\overline{I}_{n,2}+\sqrt{n}(I_{n,1}+\overline{I}_{n,3})=\mathrel{\mathop{\ordinarycolon}}\overline{Y}_{n}.$
Next, it holds for all $t=(t_{1},\ldots,t_{d})$ with $t_{j}\in\overline{\mathbb{R}}$ that
$\mathbb{P}\mathinner{\bigl(Y_{n}\leq t\bigr)}\geq\mathbb{P}\mathinner{\bigl(\overline{Y}_{n}\leq t,Y_{n}\leq\overline{Y}_{n}\bigr)}\geq\mathbb{P}\mathinner{\bigl(\overline{Y}_{n}\leq t\bigr)}+\mathbb{P}\mathinner{\bigl(Y_{n}\leq\overline{Y}_{n}\bigr)}-1\geq\mathbb{P}\mathinner{\bigl(\overline{Y}_{n}\leq t\bigr)}-\frac{4}{n}.$
Similarly,
$\mathbb{P}\mathinner{\bigl(Y_{n}\leq t\bigr)}\leq\mathbb{P}\mathinner{\bigl(\underline{Y}_{n}\leq t\bigr)}+\mathbb{P}\mathinner{\bigl(\underline{Y}_{n}>Y_{n}\bigr)}\leq\mathbb{P}\mathinner{\bigl(\underline{Y}_{n}\leq t\bigr)}+\frac{4}{n}.$
Thus,
$\sup_{t\in\overline{\mathbb{R}}^{d}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(Y_{n}\leq t\bigr)}-\mathbb{P}\mathinner{\bigl(Z\leq t\bigr)}\Bigr\rvert}$ $\leq\sup_{t\in\overline{\mathbb{R}}^{d}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(\underline{Y}_{n}\leq t\bigr)}-\mathbb{P}\mathinner{\bigl(Z\leq t\bigr)}\Bigr\rvert}$
$+\sup_{t\in\overline{\mathbb{R}}^{d}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(\overline{Y}_{n}\leq t\bigr)}-\mathbb{P}\mathinner{\bigl(Z\leq t\bigr)}\Bigr\rvert}+\frac{4}{n}.$ (C.4)
We proceed by bounding the second summand on the right-hand side of the previous display (the argument for the first summand is analogous and hence skipped). To this end, note that (C.1) and (C.2) imply that $\mathfrak{T}_{n}^{(2)}\leq b_{1}^{2}/2$ which then together with Lemma B.1 implies that $\min_{j=1,\ldots,d}\overline{\Sigma}_{\varepsilon_{n},j,j}\geq b_{1}^{2}/2$. Next, by the definition of $\overline{Y}_{n}$
$\mathbb{P}\mathinner{\bigl(\overline{Y}_{n}\leq t\bigr)}=\mathbb{P}\mathinner{\bigl(\sqrt{n}\overline{I}_{n,2}\leq t-\sqrt{n}\mathinner{(I_{n,1}+\overline{I}_{n,3})}\bigr)},\qquad t\in\overline{\mathbb{R}}^{d},$
whence, noting also that $I_{n,1}$ and $\overline{I}_{n,3}$ are non-random, Lemma B.2 applied with $\mathfrak{b}_{1}^{2}=b_{1}^{2}/2$ implies the existence of a constant $C_{3}$ depending only on $b_{1},b_{2},\lambda_{1},\lambda_{2}$, and $m$ such that
$\sup_{t\in\overline{\mathbb{R}}^{d}}\mathinner{\!\biggl\lvert\mathbb{P}\mathinner{\bigl(\overline{Y}_{n}\leq t\bigr)}-\mathbb{P}\mathinner{\biggl(\overline{Z}\leq t-\sqrt{n}\mathinner{(I_{n,1}+\overline{I}_{n,3})}\biggr)}\biggr\rvert}\leq C_{3}\mathinner{\biggl(\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr)}^{\frac{1}{4}},$
where $\overline{Z}\sim\mathsf{N}_{d}(0,\overline{\Sigma}_{\varepsilon_{n}})$. Furthermore, by the Gaussian anti-concentration inequality as stated in Theorem 1 of Chernozhukov et al. (2017b), (cf. also Lemma A.1 in Chernozhukov et al. (2017a)),
$0$ $\geq\mathbb{P}\mathinner{\bigl(\overline{Z}\leq t-\sqrt{n}\mathinner{(I_{n,1}+\overline{I}_{n,3})}\bigr)}-\mathbb{P}\mathinner{\bigl(\overline{Z}\leq t\bigr)}$
$\geq-\frac{\sqrt{n}\mathinner{(I_{n,1}+\overline{I}_{n,3})}}{\sqrt{\min_{j=1,\ldots,d}\overline{\Sigma}_{\varepsilon_{n},j,j}}}\mathinner{\bigl(\sqrt{2\log(d)}+2\bigr)}$
$\geq-C_{4}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr)}\sqrt{n\log(d)},$
the final inequality following from Lemma B.3, (C.3), $\min_{j=1,\ldots,d}\overline{\Sigma}_{\varepsilon_{n},j,j}\geq b_{1}^{2}/2$, and $C_{4}$ being a constant depending on $b_{1}$, $b_{2}$$\lambda_{1},\lambda_{2}$, and $m$ only. Thus, combining the previous two displays, there exists a constant $C_{5}$ depending on $b_{1},b_{2},\lambda_{1},\lambda_{2}$, and $m$ only such that when (C.2) is satisfied one has
$\sup_{t\in\overline{\mathbb{R}}^{d}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(\overline{Y}_{n}\leq t\bigr)}-\mathbb{P}\mathinner{\bigl(\overline{Z}\leq t\bigr)}\Bigr\rvert}\leq C_{5}\mathinner{\Biggl(\mathinner{\biggl[\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr]}^{\frac{1}{4}}+\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr]}\sqrt{n\log(d)}\Biggr)}.$
Finally, by the Gaussian-to-Gaussian comparison inequality as stated in Proposition 2.1 of Chernozhukov et al. (2022) and Lemma B.1, there exists a constant $C_{6}$, depending on $b_{1}$$b_{2}$ $\lambda_{1}$$\lambda_{2}$, and $m$ only ($C_{6}$ changes value in the second inequality below), such that
$\sup_{t\in\overline{\mathbb{R}}^{d}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(\overline{Z}\leq t\bigr)}-\mathbb{P}\mathinner{\bigl(Z\leq t\bigr)}\Bigr\rvert}$ $\leq C_{6}\mathinner{\Bigl(\log^{2}(d)\varepsilon_{n}^{1-\frac{2}{m}}\Bigr)}^{1/2}$
$\leq C_{6}\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{2}{m}}\biggr]}\Biggr)}^{1/2}.$
The previous two displays yield that there exists a constant $C_{7}$ depending only on $b_{1}$$b_{2}$$\lambda_{1},\lambda_{2}$, and $m$ such that when (C.2) is satisfied it holds that
$\sup_{t\in\overline{\mathbb{R}}^{d}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(\overline{Y}_{n}\leq t\bigr)}-\mathbb{P}\mathinner{\bigl(Z\leq t\bigr)}\Bigr\rvert}\leq C_{7}\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{2}{m}}\biggr]}\Biggr)}^{1/2}$
$+\ C_{7}\mathinner{\Biggl(\mathinner{\biggl[\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr]}^{\frac{1}{4}}+\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr]}\sqrt{n\log(d)}\Biggr)}.$
This implies the claimed bound on $\rho_{n,W}$ in (5) by using that $4/n$ in (C) is dominated by, e.g., $\mathinner{\bigl[\frac{\log(dn)}{n}\bigr]}^{1-\frac{1}{m}}\sqrt{n\log(d)}$ (if necessary adjust $C_{7}$).
If, on the other hand, (C.2) is not satisfied, then
$C_{1}\varepsilon_{n}^{1-\frac{2}{m}}\geq\frac{(b_{1}^{2}\wedge 1)}{2}\quad\iff\quad\varepsilon_{n}\geq\mathinner{\biggl[\frac{(b_{1}^{2}\wedge 1)}{2C_{1}}\biggr]}^{\frac{m}{m-2}}=\mathrel{\mathop{\ordinarycolon}}K(b_{1},b_{2},\lambda_{2},m),$
where $K=K(b_{1},b_{2},c,m)$ does not depend on $n$. Thus, by the definition of $\varepsilon_{n}$,
$\overline{\eta}_{n}\geq\frac{K}{2\lambda_{1}}\qquad\text{or}\qquad\frac{\log(dn)}{n}\geq\frac{K}{2\lambda_{2}},$
In either case, since $\rho_{n,W}\leq 1$ the bound in (5) remains valid by adjusting the constant $C$ there, if necessary.
Finally, $\rho_{n,W}\to 0$ by (B.7) and (B.8) of Lemma B.4.
Appendix D Proofs for Section 3
We start with the proof of Theorem 3.1. Before we give the proof, we first define $c_{1,n}^{\prime}$ and $c_{2,n}^{\prime}$ that show up in the proof: Recall the definition of $\varepsilon_{n}^{\prime}$ in (7). For $\lambda_{1}^{\prime}\in(1,\infty)$, define $A_{+}^{\prime}=1-\lambda_{1}^{\prime-1}\mathds{1}\mathinner{\{\overline{\eta}_{n}>0\}}$$A_{-}^{\prime}=1+\lambda_{1}^{\prime-1}\mathds{1}\mathinner{\{\overline{\eta}_{n}>0\}}$, and
$c_{1,n}^{\prime}\mathrel{\mathop{\ordinarycolon}}=-A_{+}^{\prime}W_{0}\mathinner{\bigl(-e^{-(\frac{\log(d^{2}n)}{\varepsilon_{n}^{\prime}n}+A_{+}^{\prime})/A_{+}^{\prime}}\bigr)}\in(0,A_{+}^{\prime})$ (D.1)
as well as
$c_{2,n}^{\prime}\mathrel{\mathop{\ordinarycolon}}=-A_{-}^{\prime}W_{-1}\mathinner{\bigl(-e^{-(\frac{\log(d^{2}n)}{\varepsilon_{n}^{\prime}n}+A_{-}^{\prime})/A_{-}^{\prime}}\bigr)}\in(A_{-}^{\prime},\infty);$ (D.2)
where we recall that $W_{0}$ and $W_{-1}$ denote the principal and lower branch, respectively, of Lambert’s $W$ function (cf., e.g., Corless et al. (1996)).
Proof of Theorem 3.1.
Fix $1\leq j,k\leq d$ and note that
$\tilde{\Sigma}_{n,j,k}$ $=\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\hat{a}_{j},\hat{b}_{j}}(\tilde{X}_{i,j})-\mu_{j}-(\tilde{\mu}_{n,j}-\mu_{j})\bigr]}\mathinner{\bigl[\phi_{\hat{a}_{k},\hat{b}_{k}}(\tilde{X}_{i,k})-\mu_{k}-(\tilde{\mu}_{n,k}-\mu_{k})\bigr]}$
$=\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\hat{a}_{j},\hat{b}_{j}}(\tilde{X}_{i,j})-\mu_{j}\bigr]}\mathinner{\bigl[\phi_{\hat{a}_{k},\hat{b}_{k}}(\tilde{X}_{i,k})-\mu_{k}\bigr]}-(\tilde{\mu}_{n,j}-\mu_{j})(\tilde{\mu}_{n,k}-\mu_{k})$
$=\frac{1}{n}\sum_{i=1}^{n}\phi_{\check{a}_{j},\check{b}_{j}}(\tilde{Y}_{i,j})\phi_{\check{a}_{k},\check{b}_{k}}(\tilde{Y}_{i,k})-(\tilde{\mu}_{n,j}-\mu_{j})(\tilde{\mu}_{n,k}-\mu_{k}),$
where $Y_{i,j}=X_{i,j}-\mu_{j}$$\tilde{Y}_{i,j}=\tilde{X}_{i,j}-\mu_{j}$$\check{a}_{j}=\tilde{Y}_{\lceil\varepsilon_{n}^{\prime}n\rceil,j}^{*}=\tilde{X}_{\lceil\varepsilon_{n}^{\prime}n\rceil,j}^{*}-\mu_{j}=\hat{a}_{j}-\mu_{j}$, and $\check{b}_{j}=\tilde{Y}_{\lceil(1-\varepsilon_{n}^{\prime})n\rceil,j}^{*}=\tilde{X}_{\lceil(1-\varepsilon_{n}^{\prime})n\rceil,j}^{*}-\mu_{j}=\hat{b}_{j}-\mu_{j}$.
By Theorem 2.1 in Kock and Preinerstorfer (2025) applied with $\delta=6/(d^{2}n)$ $\lambda_{i}=\lambda_{i}^{\prime}$ for $i=1,2$, there exists a constant $C_{1}$ depending only on $b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$, such that with probability at least $1-12/(d^{2}n)$
$|\tilde{\mu}_{n,j}-\mu_{j}||\tilde{\mu}_{n,k}-\mu_{k}|$ $\leq C_{1}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m\wedge 2}}\biggr)}^{2}$
$\leq C_{1}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{(m/2)\wedge 2}}\biggr)},$
the second estimate using that $\overline{\eta}_{n}$ and $\frac{\log(dn)}{n}$ are both strictly smaller than 1 by (9) (and potentially adjusting $C_{1}$).
[…]
$\mathinner{\!\Bigl\lvert\frac{1}{n}\sum_{i=1}^{n}\phi_{\check{\alpha}_{j,k},\check{\beta}_{j,k}}(\tilde{U}_{i,j,k})-\Sigma_{j,k}\Bigr\rvert}\leq C_{2}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{(m/2)\wedge 2}}\biggr)}.$
Thus, defining
$\mathfrak{S}_{n,j,k}=\mathrel{\mathop{\ordinarycolon}}\frac{1}{n}\sum_{i=1}^{n}\phi_{\check{a}_{j},\check{b}_{j}}(\tilde{Y}_{i,j})\phi_{\check{a}_{k},\check{b}_{k}}(\tilde{Y}_{i,k})-\frac{1}{n}\sum_{i=1}^{n}\phi_{\check{\alpha}_{j,k},\check{\beta}_{j,k}}(\tilde{U}_{i,j,k}),$
in order to prove the theorem, it suffices (by the union bound over $d^{2}$) to show that there exists a constant $C_{3}$ only on $b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$ such that with probability at least $1-6/(d^{2}n)$
$\mathinner{\!\bigl\lvert\mathfrak{S}_{n,j,k}\bigr\rvert}\leq C_{3}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{(m/2)\wedge 2}}\biggr)}$ (D.3)
To establish (D.3), note that with
$A_{n,j,k}\mathrel{\mathop{\ordinarycolon}}=\mathinner{\Bigl\{i\in\mathinner{\bigl\{1,\ldots,n\bigr\}}\mathrel{\mathop{\ordinarycolon}}\check{\alpha}_{j,k}\leq\tilde{U}_{i,j,k}\leq\check{\beta}_{j,k}\text{ and }\check{a}_{j}\leq\tilde{Y}_{i,j}\leq\check{b}_{j}\text{ and }\check{a}_{k}\leq\tilde{Y}_{i,k}\leq\check{b}_{k}\Bigr\}},$
one has that
$\phi_{\check{\alpha}_{j,k},\check{\beta}_{j,k}}\mathinner{(\tilde{U}_{i,j,k})}-\phi_{\check{a}_{j},\check{b}_{j}}(\tilde{Y}_{i,j})\phi_{\check{a}_{k},\check{b}_{k}}(\tilde{Y}_{i,k})=\tilde{U}_{i,j,k}-\tilde{Y}_{i,j}\tilde{Y}_{i,k}=0\qquad\text{for }i\in A_{n,j,k}.$
Hence,
$\mathfrak{S}_{n}$ $=\frac{1}{n}\sum_{i\in A_{n,j,k}^{c}}\mathinner{\bigl[\phi_{\check{\alpha}_{j,k},\check{\beta}_{j,k}}\mathinner{(\tilde{U}_{i,j,k})}-\phi_{\check{a}_{j},\check{b}_{j}}(\tilde{Y}_{i,j})\phi_{\check{a}_{k},\check{b}_{k}}(\tilde{Y}_{i,k})\bigr]},$
and for every $i\in A_{n,j,k}^{c}$
$\mathinner{\!\Bigl\lvert\phi_{\check{\alpha}_{j,k},\check{\beta}_{j,k}}\mathinner{(\tilde{U}_{i,j,k})}-\phi_{\check{a}_{j},\hat{b}_{j}}(\tilde{Y}_{i,j})\phi_{\check{a}_{k},\hat{b}_{k}}(\tilde{Y}_{i,k})\Bigr\rvert}\leq\mathinner{\bigl(|\check{\alpha}_{j,k}|\vee|\check{\beta}_{j,k}|\bigr)}+\mathinner{\bigl(|\check{a}_{j}|\vee|\check{b}_{j}|\bigr)}\mathinner{\bigl(|\check{a}_{k}|\vee|\check{b}_{k}|\bigr)}.$ (D.4)
To bound the right-hand side of (D.4), note that by Lemma G.7
$Q_{c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(U_{1,j,k})\leq\hat{\alpha}_{j,k}\leq\hat{\beta}_{j,k}\leq Q_{1-c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(U_{1,j,k})$
with probability at least $1-\frac{2}{d^{2}n}$, and where the second inequality used that $\varepsilon_{n}^{\prime}\in(0,1/2)$ by (9). Therefore, with at least this probability,
$|\check{\alpha}_{j,k}|\vee|\check{\beta}_{j,k}|\leq|Q_{c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(U_{1,j,k})|\vee|Q_{1-c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(U_{1,j,k})|$
By the same argument, it holds with probability at least $1-\frac{4}{d^{2}n}$ that
$\mathinner{\bigl(|\check{a}_{j}|\vee|\check{b}_{j}|\bigr)}\mathinner{\bigl(|\check{a}_{k}|\vee|\check{b}_{k}|\bigr)}\leq$ $\mathinner{\bigl(|Q_{c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(Y_{1,j})|\vee|Q_{1-c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(Y_{1,j})|\bigr)}$
$\cdot$ $\mathinner{\bigl(|Q_{c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(Y_{1,k})|\vee|Q_{1-c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(Y_{1,k})|\bigr)}.$
Thus, with probability at least $1-\frac{6}{d^{2}n}$ the right-hand side of (D.4) is bounded from above by
$|Q_{c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(U_{1,j,k})|\vee|Q_{1-c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(U_{1,j,k})|+$ $\mathinner{\bigl(|Q_{c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(Y_{1,j})|\vee|Q_{1-c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(Y_{1,j})|\bigr)}$
$\cdot$ $\mathinner{\bigl(|Q_{c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(Y_{1,k})|\vee|Q_{1-c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(Y_{1,k})|\bigr)}.$
Next, recall that $\mathbb{E}U_{1,j,k}=\Sigma_{j,k}$$\mathbb{E}\mathinner{\lvert U_{1,j,k}-\mathbb{E}U_{1,j,k}\rvert}^{m/2}\leq 2^{m/2}\mathbb{E}|U_{1,j,k}|^{m/2}\leq 2^{m/2}\sigma_{m}^{m}$, $\mathbb{E}Y_{1,j}=0$, and $\mathbb{E}|Y_{1,j}|^{m}\leq\sigma_{m}^{m}$ for all $1\leq j,k\leq d$. Hence, Lemma G.1 implies that the previous display is bounded from above by
$\mathinner{\biggl(|\Sigma_{j,k}|+\frac{2\sigma_{m}^{2}}{(c_{1,n}^{\prime}\varepsilon_{n}^{\prime})^{2/m}}\biggr)}+\frac{\sigma_{m}^{2}}{(c_{1,n}^{\prime}\varepsilon_{n}^{\prime})^{2/m}}\leq\frac{4\sigma_{m}^{2}}{(c_{1,n}^{\prime}\varepsilon_{n}^{\prime})^{2/m}},$
the inequality following from $|\Sigma_{j,k}|\leq\sigma_{2}^{2}\leq\sigma_{m}^{2}$. Thus, since $|A_{n,j,k}^{c}|\leq 6\varepsilon_{n}^{\prime}n$, with probability at least $1-\frac{6}{d^{2}n}$ it holds that
$\mathinner{\!\bigl\lvert\mathfrak{S}_{n,j,k}\bigr\rvert}\leq 6\varepsilon_{n}^{\prime}\cdot\frac{4\sigma_{m}^{2}}{(c_{1,n}^{\prime}\varepsilon_{n}^{\prime})^{2/m}}\leq\frac{24\sigma_{m}^{2}}{(c_{1,n}^{\prime})^{2/m}}\cdot{\varepsilon_{n}^{\prime}}^{1-\frac{2}{m}}\leq C_{3}\mathinner{\Bigl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(d^{2}n)}{n}\Bigr]}^{1-\frac{2}{m}}\Bigr)},$
where we inserted $\varepsilon_{n}^{\prime}$ from (7), used subadditivity of $z\mapsto z^{1-\frac{2}{m}}$, the lower bound on $c_{1,n}^{\prime}$ from Lemma G.6, and $C_{3}$ is the desired constant depending only on $b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$. Hence, (D.3) follows because $1-\frac{2}{m}\geq 1-\frac{1}{(m/2)\wedge 2}$ and $\frac{\log(d^{2}n)}{n}<1$ by (9).
Proof of Theorem 3.2.
Note first that writing $\mathfrak{T}_{n}^{(3)}$ for the left-hand side of (9), it follows from $\frac{\log(d^{2}n)}{n}\leq{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}$ that
$\mathfrak{T}_{n}^{(3)}\leq 2\varepsilon_{n}^{\prime}+{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}+\sqrt{{\lambda_{2}^{\prime}}^{-2}{\varepsilon_{n}^{\prime}}^{2}+4{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}^{2}}\leq C_{1}\varepsilon_{n}^{\prime},$
for a constant $C_{1}$ depending only on $\lambda_{2}^{\prime}$.
Suppose first that $C_{1}\varepsilon_{n}^{\prime}<1$ such that (9) is satisfied by the previous display. By the triangle inequality $\tilde{\rho}_{n,W}$ is bounded from above by the sum of
$\rho_{n,W}=\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(S_{n,W}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z\in H\bigr)}\Bigr\rvert}$,
where we recall that $Z\sim\mathsf{N}_{d}(0,\Sigma)$, and
$B_{n}\mathrel{\mathop{\ordinarycolon}}=\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(Z\in H\bigr)}-\mathbb{P}\mathinner{\bigl(\tilde{Z}\in H\mid\tilde{X}_{1},\ldots,\tilde{X}_{n}\bigr)}\Bigr\rvert}$.
First, $\rho_{n,W}\leq\mathfrak{A}_{n}$, where $\mathfrak{A}_{n}$ is the upper bound on $\rho_{n,W}$ in Theorem 2.1. Next, by the Gaussian-to-Gaussian comparison inequality as stated in Proposition 2.1 of Chernozhukov et al. (2022) (cf. also Proposition 2 in Chernozhukov et al. (2023a)),
$B_{n}\leq C\mathinner{\Bigl(\log^{2}(d)\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}\Bigr)}^{1/2}\leq C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\Biggr)}^{1/2},$
the last inequality holding with probability at least $1-\frac{24}{n}$ by Theorem 3.1 and $C$ being a constant depending on $b_{1},b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$ only (recall that (9) is satisfied).
If, on the other hand, $C_{1}\varepsilon_{n}^{\prime}\geq 1$ then by the definition of $\varepsilon_{n}^{\prime}$ in (7)
$\overline{\eta}_{n}\geq\frac{1}{2C_{1}\lambda_{1}^{\prime}}\qquad\text{or}\qquad\frac{\log(dn)}{n}=\frac{\log([dn]^{2})/2}{n}\geq\frac{\log(d^{2}n)}{2n}\geq\frac{1}{4C_{1}\lambda_{2}^{\prime}}.$
In either case, the upper bound on $\hat{\rho}_{n,W}$ remains valid by suitably adjusting $C$ (if needed).
That $\tilde{\rho}_{n,W}\to 0$ in probability follows from (B.7)–(B.9) of Lemma B.4. ∎
Appendix E Proofs for Section 4
The following “two-sided” Gaussian anti-concentration inequality is a simple consequence of the “one-sided” one stated in Theorem 1 of Chernozhukov et al. (2017b), (cf. also Lemma A.1 in Chernozhukov et al. (2017a)), but as we could not pinpoint it in the literature we state it here for completeness. For all $u,v$ in $\overline{\mathbb{R}}^{d}$, we define the set $u,v]=\mathinner{\bigl\{x\in\mathbb{R}^{d}\mathrel{\mathop{\ordinarycolon}}u_{j}\leq x_{j}\leq v_{j}\text{ for all }j=1,\ldots,d\bigr\}}$, which may be empty. Recall the notational conventions introduced after (A.10).
Lemma E.1.
Let $Z$ in $\mathbb{R}^{d}$ with $d\geq 1$ be such that $Z\sim~\mathsf{N}_{d}(0,\Sigma)$ with $\Sigma_{j,j}\geq\underline{\sigma}^{2}$ for all $j=1,\ldots,d$ and some $\underline{\sigma}^{2}>0$. Then, for all real numbers $\delta_{1}$ and $\delta_{2}$ and all $a$ and $b$ in $\overline{\mathbb{R}}^{d}$, it holds that
$\mathbb{P}\mathinner{\bigl(Z\in[a+\delta_{1},b+\delta_{2}]\bigr)}\leq\mathbb{P}\mathinner{\bigl(Z\in[a,b]\bigr)}+\frac{\overline{\delta}}{\underline{\sigma}}\mathinner{\bigl(\sqrt{2\log(d)}+4\bigr)},$ (E.1)
where $\overline{\delta}=|\delta_{1}|\vee|\delta_{2}|$, and
$\mathbb{P}\mathinner{\bigl(Z\in[a+\delta_{1},b+\delta_{2}]\bigr)}\geq\mathbb{P}\mathinner{\bigl(Z\in[a,b]\bigr)}-\frac{\overline{\delta}}{\underline{\sigma}}\mathinner{\bigl(\sqrt{2\log(d)}+4\bigr)}.$ (E.2)
Proof.
Consider first (E.1). Clearly,
$\mathbb{P}\mathinner{\bigl(Z\in[a+\delta_{1},b+\delta_{2}]\bigr)}\leq\mathbb{P}\mathinner{\bigl(Z\in[a-\overline{\delta},b+\overline{\delta}]\bigr)}=\mathbb{P}\mathinner{\bigl((Z^{\prime},-Z^{\prime})^{\prime}\leq(b^{\prime}+\overline{\delta},-a^{\prime}+\overline{\delta})^{\prime}\bigr)}.$
By Theorem 1 of Chernozhukov et al. (2017b), which trivially remains valid for $y$ there taking values in $\overline{\mathbb{R}}^{d}$, the far right-hand side of the previous display is bounded from above by the sum of $\mathbb{P}\mathinner{(Z\in[a,b])}$ and
$\frac{\overline{\delta}}{\underline{\sigma}}\mathinner{\bigl(\sqrt{2\log(2d)}+2\bigr)}=\frac{\overline{\delta}}{\underline{\sigma}}\mathinner{\bigl(\sqrt{2(\log(2)+\log(d))}+2\bigr)}\leq\frac{\overline{\delta}}{\underline{\sigma}}\mathinner{\bigl(\sqrt{2\log(d)}+2+\sqrt{2(\log(2)}\bigr)},$
which yields the desired result. To prove (E.2), note that
$\mathbb{P}\mathinner{\bigl(Z\in[a+\delta_{1},b+\delta_{2}]\bigr)}\geq\mathbb{P}\mathinner{\bigl(Z\in[a+\overline{\delta},b-\overline{\delta}]\bigr)}=\mathbb{P}\mathinner{\bigl((Z^{\prime},-Z^{\prime})\leq(b^{\prime}-\overline{\delta},-a^{\prime}-\overline{\delta})^{\prime}\bigr)},$
which, by Theorem 1 of Chernozhukov et al. (2017b), is bounded from below by $\mathbb{P}\mathinner{(Z\in[a,b])}$ minus the left-hand side of the penultimate display, implying (E.2). ∎
We will use the following notation in the proof of Theorem 4.1. For $x\in\mathbb{R}^{d}$, let $\mathinner{\lVert x\rVert}_{\infty}=\max_{j=1,\ldots,d}|x_{j}|$. For any $A\subseteq\mathbb{R}^{d}$ and $\zeta>0$, let
$A^{\zeta,\infty}=\mathinner{\bigl\{x\in\mathbb{R}^{d}\mathrel{\mathop{\ordinarycolon}}\inf_{y\in A}\|x-y\|_{\infty}\leq\zeta\bigr\}}.$
Furthermore,
$A^{-\zeta,\infty}=\mathinner{\bigl\{x\in\mathbb{R}^{d}\mathrel{\mathop{\ordinarycolon}}\mathcal{B}_{\infty}(x,\zeta)\subseteq A\bigr\}}\quad\text{ where }\quad\mathcal{B}_{\infty}(x,\zeta)=\mathinner{\bigl\{y\in\mathbb{R}^{d}\mathrel{\mathop{\ordinarycolon}}\|y-x\|_{\infty}\leq\zeta\bigr\}}.$
Proof of Theorem 4.1.
We first establish (4.1). Writing $\mathfrak{T}_{n}^{(3)}$ for the left-hand side of (9), it follows from $\frac{\log(d^{2}n)}{n}\leq{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}$ that
$\mathfrak{T}_{n}^{(3)}\leq 2\varepsilon_{n}^{\prime}+{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}+\sqrt{{\lambda_{2}^{\prime}}^{-2}{\varepsilon_{n}^{\prime}}^{2}+4{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}^{2}}\leq K\varepsilon_{n}^{\prime},$
for a constant $K$ depending only on $\lambda_{2}^{\prime}$. Assume that
$K\varepsilon_{n}^{\prime}<1\qquad\text{and}\qquad C_{1}\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\leq\frac{b_{1}^{2}}{2},$ (E.3)
where $C_{1}$ is the constant from Theorem 3.1 depending only on $b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$. Note that $K\varepsilon_{n}^{\prime}<1$ implies, by the penultimate display, that (9) is satisfied which is used when invoking Theorem 3.1 below. If either of the conditions in (E.3) is violated, we conclude (4.1) with similar arguments as those employed in the end of the proofs of Theorems 2.1 and 3.2, respectively.
Recall the definition of $S_{n,W,S}$ in (11), let $T_{n}\in\mathbb{R}^{d}$ have entries
$T_{n,j}=\frac{1}{\sqrt{n}\sigma_{2,j}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}(\tilde{X}_{i,j})-\mu_{j}\bigr]},\qquad j=1,\ldots,d,$
and, observe that (grant the quotients are well-defined)
$A_{n}\mathrel{\mathop{\ordinarycolon}}=\mathinner{\!\bigl\lVert S_{n,W,S}-T_{n}\bigr\rVert}_{\infty}\leq\max_{j=1,\ldots,d}\mathinner{\!\biggl\lvert\frac{1}{\tilde{\sigma}_{n,j}}-\frac{1}{\sigma_{2,j}}\biggr\rvert}\cdot\max_{j=1,\ldots,d}\mathinner{\!\biggl\lvert\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}(\tilde{X}_{i,j})-\mu_{j}\bigr]}\biggr\rvert}$. (E.4)
[…]
$\max_{j=1,\ldots,d}\mathinner{\!\biggl\lvert\frac{1}{\tilde{\sigma}_{n,j}}-\frac{1}{\sigma_{2,j}}\biggr\rvert}=\max_{j=1,\ldots,d}\frac{|\tilde{\sigma}_{n,j}-\sigma_{2,j}|}{\tilde{\sigma}_{n,j}\sigma_{2,j}}\leq C_{2}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr)}.$ (E.5)
Furthermore, as $\sigma_{2}\leq\sigma_{m}\leq b_{2}$, the union bound and $\mathbb{P}\mathinner{(|\bm{z}|>t)}\leq 2\exp(-t^{2}/2)$ for $t\geq 0$ and $\bm{z}\sim\mathsf{N}_{1}(0,1)$ yields for $Z\sim\mathsf{N}_{d}(0,\Sigma)$ that
$\mathbb{P}\mathinner{\left(\max_{j=1,\ldots,d}|Z_{j}|>b_{2}\sqrt{2\log(2dn)}\right)}\leq\frac{1}{n}.$
Thus, by Theorem 2.1, and writing
$r_{n}\mathrel{\mathop{\ordinarycolon}}$= $\mathinner{\biggl[\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr]}^{\frac{1}{4}}+\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr]}\sqrt{n\log(d)}$
$+\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{2}{m}}\biggr]}\Biggr)}^{1/2},$
it follows that there exists a constant $K_{1}$ depending only on $b_{1},b_{2},\lambda_{1},\lambda_{2},$ and $m$ such that
$\mathbb{P}\mathinner{\biggl(\max_{j=1,\ldots,d}\mathinner{\!\biggl\lvert\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}(\tilde{X}_{i,j})-\mu_{j}\bigr]}\biggr\rvert}>b_{2}\sqrt{2\log(2dn)}\biggr)}\leq K_{1}r_{n}+\frac{1}{n}\leq K_{1}r_{n},$ (E.6)
where the value of $K_{1}$ is suitably adjusted to justify the second inequality. Hence, by (E.4)–(E.6) there exists a constant $C_{3}$ depending only on $b_{1},b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$ such that
$\mathbb{P}\mathinner{\biggl(A_{n}\leq C_{3}\sqrt{\log(dn)}\mathinner{\Bigl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\Bigr]}\biggr)}\geq 1-\frac{24}{n}-K_{1}r_{n}\geq 1-Kr_{n},$ (E.7)
where $K$ depends only on $b_{1},b_{2},\lambda_{1},\lambda_{2}$, and $m$. Thus, writing
$\overline{A}_{n}\mathrel{\mathop{\ordinarycolon}}=C_{3}\sqrt{\log(dn)}\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]},$
it holds for all $H\in\mathcal{H}$ that
$\mathinner{\bigl\{S_{n,W,S}\in H\bigr\}}\subseteq\mathinner{\bigl\{T_{n}\in H^{\overline{A}_{n},\infty}\bigr\}}\bigcup\mathinner{\bigl\{A_{n}>\overline{A}_{n}\bigr\}}.$ (E.8)
Writing $Y_{i}=D^{-1}X_{i}$$\tilde{Y}_{i}=D^{-1}\tilde{X}_{i}$ as well as $\check{\alpha}_{j}=\tilde{Y}_{\lceil\varepsilon_{n}n\rceil,j}^{*}$ and $\check{\beta}_{j}=\tilde{Y}_{\lceil(1-\varepsilon_{n})n\rceil,j}^{*}$, note that
$T_{n,j}=\frac{1}{\sqrt{n}\sigma_{2,j}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}(\tilde{X}_{i,j})-\mu_{j}\bigr]}=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\check{\alpha}_{j},\check{\beta}_{j}}(\tilde{Y}_{i,j})-\mathbb{E}Y_{1,j}\bigr]},\qquad\text{for }j=1,\ldots,d,$
such that by Theorem 2.1 and the covariance matrix of $Y_{1}$ being $\Sigma_{0}=D^{-1}\Sigma D^{-1}$, there exists a constant $C$ depending only on $b_{1},b_{2},\lambda_{1},\lambda_{2}$, and $m$ such that (using that $H^{\overline{A}_{n},\infty}\in\mathcal{H}$)
$\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(T_{n}\in H^{\overline{A}_{n},\infty}\bigr)}-\mathbb{P}\mathinner{\bigl(Z^{\prime}\in H^{\overline{A}_{n},\infty}\bigr)}\Bigr\rvert}\leq Cr_{n}.$
Furthermore, by Lemma E.1 (applied with $\delta_{1}=-\overline{A}_{n}$ and $\delta_{2}=\overline{A}_{n}$) and $Z^{\prime}\sim\mathsf{N}_{d}(0,\Sigma_{0})$ where $\Sigma_{0,j,j}=1$ for all $j=1,\ldots,d$,
$0\leq\mathbb{P}\mathinner{\bigl(Z^{\prime}\in H^{\overline{A}_{n},\infty}\bigr)}-\mathbb{P}\mathinner{\bigl(Z^{\prime}\in H\bigr)}\leq\overline{A}_{n}\mathinner{\bigl(\sqrt{2\log(d)}+4\bigr)}\leq 7\overline{A}_{n}\sqrt{\log(d)},$
the last inequality following from $d\geq 2$. Using this in (E.8) along with (E.7) yields that
$\mathbb{P}\mathinner{\bigl(S_{n,W,S}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z_{j}^{\prime}\in H\bigr)}\leq(C+K)r_{n}+7\overline{A}_{n}\sqrt{\log(d)}.$
Finally, since
$\mathinner{\bigl\{T_{n}\in H^{-\overline{A}_{n},\infty}\bigr\}}\subseteq\mathinner{\bigl\{S_{n,W,S}\in H\bigr\}}\bigcup\mathinner{\bigl\{A_{n}>\overline{A}_{n}\bigr\}},$
the same arguments as those following (E.8) lead to
$(C+K)r_{n}-7\overline{A}_{n}\sqrt{\log(d)}\leq\mathbb{P}\mathinner{\bigl(S_{n,W,S}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z^{\prime}\in H\bigr)},$
which yields (4.1) under the assumption of (E.3), upon adjusting constants as the upper and lower bounds just obtained do not depend on $H$.
Finally, since $r_{n}\to 0$ by (B.7) and (B.8) as well as $\overline{A}_{n}\sqrt{\log(d)}\to 0$ by (B.10) of Lemma B.4 it follows that $\rho_{n,W,S}\to 0$. ∎
Proof of Theorem 4.2.
The proof is similar to that of Theorem 3.2, but we include it here for completeness. By the triangle inequality $\tilde{\rho}_{n,W,S}$ is bounded from above by the sum of
$\rho_{n,W,S}=\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(S_{n,W,S}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z^{\prime}\in H\bigr)}\Bigr\rvert}$,
where we recall that $Z^{\prime}\sim\mathsf{N}_{d}(0,\Sigma_{0})$, and
$B_{n}\mathrel{\mathop{\ordinarycolon}}=\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(Z^{\prime}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(\hat{Z}^{\prime}\in H\mid\tilde{X}_{1},\ldots,\tilde{X}_{n}\bigr)}\Bigr\rvert}$.
First, $\rho_{n,W,S}\leq\mathfrak{B}_{n}$, where $\mathfrak{B}_{n}$ is the upper bound on $\rho_{n,W,S}$ in Theorem 4.1. Next, by the Gaussian-to-Gaussian comparison inequality as stated in Proposition 2.1 of Chernozhukov et al. (2022) (cf. also Proposition 2 in Chernozhukov et al. (2023a)),
$B_{n}\leq C\mathinner{\Bigl(\log^{2}(d)\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,0,j,k}-\Sigma_{0,j,k}\bigr\rvert}\Bigr)}^{1/2},$
for an absolute constant $C$ (since $\Sigma_{0,j,j}=1$ for all $j=1,\ldots,d$).
Arguing as in the opening paragraph of the proof of Theorem 4.1, it suffices to consider the case of (E.3) being satisfied.
Let $C_{1}$ be the constant from Theorem 3.1 depending only on $b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$. By that theorem (note that (9) is ensured to be true by (E.3) so that Theorem 3.1 can be employed) there exists a set $E_{n}$ of probability at least $1-\frac{24}{n}$ on which
$\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}\leq C_{1}\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\leq\frac{b_{1}^{2}}{2}.$
Therefore, on $E_{n}$$\min_{j=1,\ldots,d}\tilde{\sigma}_{n,j}^{2}\geq b_{1}^{2}/2$ because by assumption $\min_{j=1,\ldots,d}\sigma_{2,j}^{2}\geq b_{1}^{2}$. Furthermore, note that for all $1\leq j,k\leq d$
$\tilde{\Sigma}_{n,0,j,k}-\Sigma_{0,j,k}=\frac{\tilde{\Sigma}_{n,j,k}}{\tilde{\sigma}_{n,j}\tilde{\sigma}_{n,j}}-\frac{\Sigma_{j,k}}{\sigma_{2,j}\sigma_{2,k}}=\frac{(\tilde{\Sigma}_{n,j,k}-\Sigma_{j,k})\sigma_{2,j}\sigma_{2,k}+\Sigma_{j,k}(\sigma_{2,j}\sigma_{2,k}-\tilde{\sigma}_{n,j}\tilde{\sigma}_{n,k})}{\tilde{\sigma}_{n,j}\tilde{\sigma}_{n,k}\sigma_{2,j}\sigma_{2,k}},$
which is well-defined on $E_{n}$. Thus, by the mean-value theorem, there exists a constant $C_{2}$ depending only on $b_{1},b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$ such that on $E_{n}$.
$\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,0,j,k}-\Sigma_{0,j,k}\bigr\rvert}\leq C_{2}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr)}.$
Hence, with probability at least $1-\frac{24}{n}$
$B_{n}\leq C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\Biggr)}^{1/2}$
for a constant $C$ depending only on $b_{1},b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$ which implies the bound on $\tilde{\rho}_{n,W,S}$ in (13) in case of (E.3).
Finally, $\tilde{\rho}_{n,W,S}\to 0$ by (B.7)–(B.10) of Lemma B.4. ∎
Appendix F Proofs for Section 5
Proof of Theorem 5.1.
Since $\log(dn)/n\leq\lambda_{2}^{-1}\varepsilon_{n}$, one has that $\mathfrak{T}_{n}^{(1)}$ in (9) is bounded as
$\mathfrak{T}_{n}^{(1)}\leq 2\varepsilon_{n}+\lambda_{2}^{-1}\varepsilon_{n}+\sqrt{\lambda_{2}^{-2}\varepsilon_{n}^{2}+4\lambda_{2}^{-1}\varepsilon_{n}^{2}}\leq C_{1}\varepsilon_{n},$
for a constant $C_{1}$ depending only on $\lambda_{2}$. In the sequel we impose that $C_{1}\varepsilon_{n}<1/2$ such that by the previous display (A.5) is satisfied allowing us to invoke Lemma G.3 below. In case $C_{1}\varepsilon_{n}\geq 1/2$, (15) trivially holds by the same arguments as in the end of the proof of Theorem 2.1.
Fix $j\in\mathinner{\{1,\ldots,d\}}$. Since $\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}(\tilde{X}_{i,j})=\tilde{X}_{i,j}$ for $i\in I_{n}=\mathinner{\bigl\{\lceil\varepsilon_{n}n\rceil,\ldots,\lceil(1-\varepsilon_{n})n\rceil\bigr\}}$
$A_{n,j}\mathrel{\mathop{\ordinarycolon}}$ $=\mathinner{\!\bigl\lvert S_{n,T,j}-S_{n,W,j}\bigr\rvert}$
$=\sqrt{n}\mathinner{\!\Biggl\lvert\frac{1}{|I_{n}|}\sum_{i=\lceil\varepsilon_{n}n\rceil}^{\lceil(1-\varepsilon_{n})n\rceil}\mathinner{\bigl[\tilde{X}_{i,j}^{*}-\mu_{j}\bigr]}-\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}(\tilde{X}_{i,j}^{*})-\mu_{j}\bigr]}\Biggr\rvert}$
$=\sqrt{n}\mathinner{\!\Biggl\lvert\frac{\lfloor\varepsilon_{n}n\rfloor+\lceil\varepsilon_{n}n\rceil-1}{n|I_{n}|}\sum_{i=\lceil\varepsilon_{n}n\rceil}^{\lceil(1-\varepsilon_{n})n\rceil}\mathinner{\bigl[\tilde{X}_{i,j}^{*}-\mu_{j}\bigr]}-\frac{\lceil\varepsilon_{n}n\rceil-1}{n}\mathinner{\bigl[\hat{\alpha}_{j}-\mu_{j}\bigr]}-\frac{\lfloor\varepsilon_{n}n\rfloor}{n}\mathinner{\bigl[\hat{\beta}_{j}-\mu_{j}\bigr]}\Biggr\rvert}$
$\leq\sqrt{n}\mathinner{\Biggl(\frac{2\lceil{\varepsilon_{n}n}\rceil}{n|I_{n}|}\sum_{i=\lceil\varepsilon_{n}n\rceil}^{\lceil(1-\varepsilon_{n})n\rceil}\mathinner{\!\bigl\lvert\tilde{X}_{i,j}^{*}-\mu_{j}\bigr\rvert}+\frac{\lceil\varepsilon_{n}n\rceil}{n}\mathinner{\!\bigl\lvert\hat{\alpha}_{j}-\mu_{j}\bigr\rvert}+\frac{\lceil\varepsilon_{n}n\rceil}{n}\mathinner{\!\bigl\lvert\hat{\beta}_{j}-\mu_{j}\bigr\rvert}\Biggr)}$.
Next, since
$\frac{2\lceil{\varepsilon_{n}n}\rceil}{n|I_{n}|}\sum_{i=\lceil\varepsilon_{n}n\rceil}^{\lceil(1-\varepsilon_{n})n\rceil}\mathinner{\!\bigl\lvert\tilde{X}_{i,j}^{*}-\mu_{j}\bigr\rvert}$ $\leq\frac{2\lceil{\varepsilon_{n}n}\rceil}{n}\max_{i=\lceil\varepsilon_{n}n\rceil,\ldots,\lceil(1-\varepsilon_{n})n\rceil}\mathinner{\!\bigl\lvert\tilde{X}_{i,j}^{*}-\mu_{j}\bigr\rvert}$
$=\frac{2\lceil{\varepsilon_{n}n}\rceil}{n}\mathinner{\bigl[|\hat{\alpha}_{j}-\mu_{j}|\vee|\hat{\beta}_{j}-\mu_{j}|\bigr]},$
it suffices to bound $|\hat{\alpha}_{j}-\mu_{j}|\vee|\hat{\beta}_{j}-\mu_{j}|$ from above. To this end, by Lemma G.3 and $\hat{\alpha}_{j}-\mu_{j}\leq\hat{\beta}_{j}-\mu_{j}$ ((A.5) implies $\varepsilon_{n}<1/4$ such that $\hat{\alpha}_{j}=\tilde{X}_{\lceil{\varepsilon_{n}n}\rceil,j}^{*}\leq\tilde{X}_{\lceil(1-\varepsilon_{n})n\rceil,j}^{*}=\hat{\beta}_{j}$) it holds with probability at least $1-\frac{2}{dn}$ that
$|\hat{\alpha}_{j}-\mu_{j}|\vee|\hat{\beta}_{j}-\mu_{j}|\leq|Q_{c_{1,n}\varepsilon_{n}}(X_{1,j})-\mu_{j}|\vee|Q_{1-c_{1,n}\varepsilon_{n}}(X_{1,j})-\mu_{j}|\leq\frac{\sigma_{m}}{(c_{1,n}\varepsilon_{n})^{1/m}},$
where the last estimate is by Lemma G.1. Therefore, by Lemma G.2 there exists a constant $C$ depending only on $b_{2},\lambda_{1},\lambda_{2}$, and $m$ such that with probability at least $1-\frac{2}{dn}$,
$A_{n,j}\leq\frac{4\sqrt{n}\lceil{\varepsilon_{n}n}\rceil\sigma_{m}}{n(c_{1,n}\varepsilon_{n})^{1/m}}\leq C\sqrt{n}\varepsilon_{n}^{1-\frac{1}{m}}\leq C\sqrt{n}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr)}=\mathrel{\mathop{\ordinarycolon}}\overline{A}_{n},$
with $C$ potentially changing values in the last inequality and subadditivity of $z\mapsto z^{1-\frac{1}{m}}$ was used along with the definition of $\varepsilon_{n}$ in (4). Therefore, since the right-hand side does not depend on $j$, it follows by the union bound over $j=1,\ldots,d$ that with probability at least $1-\frac{2}{n}$
$A_{n}\mathrel{\mathop{\ordinarycolon}}=\max_{j=1,\ldots,d}A_{n,j}\leq\overline{A}_{n}.$
Next, observe that for all $H\in\mathcal{H}$ (recalling also the notation introduced prior to the proof of Theorem 4.1)
$\mathinner{\bigl\{S_{n,T}\in H\bigr\}}\subseteq\mathinner{\bigl\{S_{n,W}\in H^{\overline{A}_{n},\infty}\bigr\}}\bigcup\mathinner{\bigl\{A_{n}>\overline{A}_{n}\bigr\}}.$
[…]
$\mathbb{P}\mathinner{\bigl(S_{n,W}\in H^{\overline{A}_{n},\infty}\bigr)}-\mathbb{P}\mathinner{\bigl(Z\in H\bigr)}\leq\mathfrak{A}_{n}+\frac{\overline{A}_{n}}{b_{1}}\mathinner{\bigl(\sqrt{2\log(d)}+4\bigr)}\leq\mathfrak{A}_{n}+C\overline{A}_{n}\sqrt{\log(d)},$
for a constant $C$ depending only on $b_{1}$. Thus, since  $\mathbb{P}\mathinner{\bigl(A_{n}>\overline{A}_{n}\bigr)}\leq\frac{2}{n}$, we conclude that
$\mathbb{P}\mathinner{\bigl(S_{n,T}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z\in H\bigr)}\leq\mathfrak{A}_{n}+C\overline{A}_{n}\sqrt{\log(d)}+\frac{2}{n}.$ (F.1)
Since also
$\mathinner{\bigl\{S_{n,W}\in H^{-\overline{A}_{n},\infty}\bigr\}}\subseteq\mathinner{\bigl\{S_{n,T}\in H\bigr\}}\bigcup\mathinner{\bigl\{A_{n}>\overline{A}_{n}\bigr\}},$
an identical argument shows that
$\mathfrak{A}_{n}-C\overline{A}_{n}\sqrt{\log(d)}-\frac{2}{n}\leq\mathbb{P}\mathinner{\bigl(S_{n,T}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z\in H\bigr)},$ (F.2)
which, together with the penultimate display, (F.1)–(F.2) not depending on $H\in\mathcal{H}$, and dominating $2/n$ by $\overline{A}_{n}\sqrt{\log(d)}$ implies the bound in (15). That $\rho_{n,T}\to 0$ follows from (B.7) and (B.8) of Lemma B.4. ∎
Proof of Theorem 5.2.
The proof is almost identical to that of Theorem 3.2, but is included for completeness. Note first that writing $\mathfrak{T}_{n}^{(3)}$ for the left-hand side of (9), it follows from $\frac{\log(d^{2}n)}{n}\leq{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}$ that
$\mathfrak{T}_{n}^{(3)}\leq 2\varepsilon_{n}^{\prime}+{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}+\sqrt{{\lambda_{2}^{\prime}}^{-2}{\varepsilon_{n}^{\prime}}^{2}+4{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}^{2}}\leq C_{1}\varepsilon_{n}^{\prime},$
for a constant $C_{1}$ depending only on $\lambda_{2}^{\prime}$.
Suppose first that $C_{1}\varepsilon_{n}^{\prime}<1$ such that (9) is satisfied by the previous display. By the triangle inequality $\tilde{\rho}_{n,T}$ is bounded from above by the sum of
$\rho_{n,T}=\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(S_{n,T}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z\in H\bigr)}\Bigr\rvert}$,
where we recall that $Z\sim\mathsf{N}_{d}(0,\Sigma)$, and
$B_{n}\mathrel{\mathop{\ordinarycolon}}=\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(Z\in H\bigr)}-\mathbb{P}\mathinner{\bigl(\tilde{Z}\in H\mid\tilde{X}_{1},\ldots,\tilde{X}_{n}\bigr)}\Bigr\rvert}$.
First, $\rho_{n,T}\leq\mathfrak{C}_{n}$, where $\mathfrak{C}_{n}$ is the upper bound on $\rho_{n,T}$ in Theorem 5.1. Next, by the Gaussian-to-Gaussian comparison inequality as stated in Proposition 2.1 of Chernozhukov et al. (2022) (cf. also Proposition 2 in Chernozhukov et al. (2023a)),
$B_{n}\leq C\mathinner{\Bigl(\log^{2}(d)\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}\Bigr)}^{1/2}\leq C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\Biggr)}^{1/2},$
the last inequality holding with probability at least $1-\frac{24}{n}$ by Theorem 3.1 and $C$ being a constant depending on $b_{1},b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$ and $m$ only.
If, on the other hand, $C_{1}\varepsilon_{n}^{\prime}\geq 1$ then by the definition of $\varepsilon_{n}^{\prime}$ in (7)
$\overline{\eta}_{n}\geq\frac{1}{2C_{1}\lambda_{1}^{\prime}}\qquad\text{or}\qquad\frac{\log(dn)}{n}=\frac{\log([dn]^{2})/2}{n}\geq\frac{\log(d^{2}n)}{2n}\geq\frac{1}{4C_{1}\lambda_{2}^{\prime}}.$
In either case, the upper bound on $\tilde{\rho}_{n,T}$ remains valid by suitably adjusting $C$ (if needed).
That $\tilde{\rho}_{n,T}\to 0$ in probability follows from (B.7)–(B.9) of Lemma B.4. ∎
Appendix G Auxiliary lemmas
This section gathers some auxiliary lemmas, the proofs of which largely follow from related results in Kock and Preinerstorfer (2025).
The following standard result is Lemma C.1 from Kock and Preinerstorfer (2025). It bounds the difference between the mean and quantile of a distribution of a random variable $Z$ (which is not necessarily continuous).
Lemma G.1.
Let $Z$ satisfy $\sigma_{m}^{m}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}|Z-\mathbb{E}Z|^{m}\in[0,\infty)$ for some $m\in[1,\infty)$. Then, for all $p\in(0,1)$,
$\mathbb{E}Z-\frac{\sigma_{m}}{p^{1/m}}\leq Q_{p}(Z)\leq\mathbb{E}Z+\frac{\sigma_{m}}{(1-p)^{1/m}}.$ (G.1)
Lemma G.2.
For $c_{1,n}$ as defined in (A.1), $c_{2,n}$ as defined in (A.2), and $\varepsilon_{n}$ as defined in (4) and satisfying (A.5), it holds that
$0<(1-\lambda_{1}^{-1})\exp\mathinner{\Bigl({-\frac{1}{\lambda_{2}(1-\lambda_{1}^{-1})}-1}\Bigr)}\leq\$$c_{1,n}<1,$
$1<\$$c_{2,n}\leq 2+\lambda_{2}^{-1}+\sqrt{\lambda_{2}^{-2}+4\lambda_{2}^{-1}},$
and
$0<\varepsilon_{n}\min(c_{1,n},c_{2,n})$ $\leq\varepsilon_{n}(c_{1,n}+c_{2,n})$
$\leq 2\varepsilon_{n}+\frac{\log(dn)}{n}+\sqrt{\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{2}+4\frac{\log(dn)}{n}\varepsilon_{n}}.$ (G.2)
Proof.
Apply Lemma B.3 in Kock and Preinerstorfer (2025) with $\delta=\frac{6}{dn}$ (recall that $dn>6$ is assumed throughout; cf. the sentence right before Theorem 2.1), $\epsilon=\varepsilon_{n}$, and $\eta=\overline{\eta}_{n}$. To this end, note that $c_{1}$, and $c_{2}$ there equal $c_{1,n}$ and $c_{2,n}$, respectively, for $\delta=\frac{6}{dn}$, that $\varepsilon_{n}\geq\lambda_{2}\log(dn)/n$ by definition, and that $\varepsilon_{n}<1/4$ follows from (A.5). ∎
The following lemma shows that for $\varepsilon_{n}$ as defined in (4), the lower and upper $\varepsilon_{n}$ order statistics of the contaminated data are close to related population quantiles of the uncontaminated data.
Lemma G.3.
Fix $j\in\mathinner{\{1,\ldots,d\}}$$n\in\mathbb{N}$, let Assumption 2.1 be satisfied, and $\varepsilon_{n}$ be as defined in (4) satisfying (A.5). Then, for $c_{1,n}$ as in (A.1) and $c_{2,n}$ as in (A.2), we have
$0<\varepsilon_{n}\min(c_{1,n},c_{2,n})\leq\varepsilon_{n}(c_{1,n}+c_{2,n})<1/2,$
and each of (G.3)–(G.6) below holds with probability at least $1-\frac{1}{dn}$:
$\tilde{X}_{\lceil\varepsilon_{n}n\rceil,j}^{*}$ $\geq$ $Q_{c_{1,n}\varepsilon_{n}}(X_{1,j});$ (G.3)
$\tilde{X}_{\lceil(1-\varepsilon_{n})n\rceil,j}^{*}$ $\geq$ $Q_{1-c_{2,n}\varepsilon_{n}}(X_{1,j});$ (G.4)
$\tilde{X}_{\lfloor\varepsilon_{n}n\rfloor+1,j}^{*}$ $\leq$ $Q_{c_{2,n}\varepsilon_{n}}(X_{1,j});$ (G.5)
$\tilde{X}_{\lfloor(1-\varepsilon_{n})n\rfloor+1,j}^{*}$ $\leq$ $Q_{1-c_{1,n}\varepsilon_{n}}(X_{1,j}).$ (G.6)
Proof.
The first statement in the present lemma follows from (A.5) along with (G.2) of Lemma G.2. For the remaining statements, apply Lemma B.5 in Kock and Preinerstorfer (2025) with $\delta=\frac{6}{dn}$ (recall that $dn>6$ is assumed throughout; cf. the sentence right before Theorem 2.1), $\epsilon=\varepsilon_{n}$, and $\eta=\overline{\eta}_{n}$. To this end, note that
  1. 1.
    $c_{1}$, and $c_{2}$ there equal $c_{1,n}$ and $c_{2,n}$, respectively, for $\delta=\frac{6}{dn}$,
  2. 2.
    that $\varepsilon_{n}\geq\lambda_{1}\overline{\eta}_{n}$ by definition and $\varepsilon_{n}c_{2,n}<1/2$ (which has already been verified),
  3. 3.
    $\varepsilon_{n}<1/4$ follows from (A.5),
and that our Assumption 2.1 implies the remaining assumptions there. ∎
Lemma G.4.
There exists a positive constant $C$ depending only on $\lambda_{1},\lambda_{2}$, and $m$, such that if $\varepsilon_{n}$ is as in (4) satisfying (A.5) and Assumption 2.1 is satisfied, then for all $j=1,\ldots,d$,
$\mathinner{\!\biggl\lvert\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(\tilde{X}_{i,j})-\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{i,j})\bigr]}\biggr\rvert}\leq C\sigma_{m}\overline{\eta}_{n}^{1-\frac{1}{m}}$ (G.7)
and
$\mathinner{\!\biggl\lvert\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(\tilde{X}_{i,j})-\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})\bigr]}\biggr\rvert}\leq C\sigma_{m}\overline{\eta}_{n}^{1-\frac{1}{m}}.$ (G.8)
Proof.
Fix $j\in\mathinner{\{1,\ldots,d\}}$ and recall the definitions of $\underline{\alpha}_{j},\overline{\alpha}_{j},\underline{\beta}_{j}$ and $\overline{\beta}_{j}$ from (A.3) and (A.4). To establish (G.7), apply Lemma C.2 in Kock and Preinerstorfer (2025) to $X_{1,j}$ with $\eta=\overline{\eta}_{n}$$s_{1}=c_{1,n}\varepsilon_{n}$, and $s_{2}=1-c_{2,n}\varepsilon_{n}$, recalling that $\sigma_{m,j}\leq\sigma_{m}$ and using that $\varepsilon_{n}>\lambda_{1}\overline{\eta}_{n}>\overline{\eta}_{n}$ as well as the lower bounds on $c_{1,n}$ and $c_{2,n}$ from Lemma G.2, noting that our Assumption 2.1 implies the assumptions there. The requirement $s_{1}<s_{2}$ in Lemma C.2 in Kock and Preinerstorfer (2025) follows by (A.5) and Lemma G.2. (G.8) follows in a similar fashion using $s_{1}=c_{2,n}\varepsilon_{n}$ and $s_{2}=1-c_{1,n}\varepsilon_{n}$ instead. ∎
Lemma G.5.
There exists a positive constant $C$ depending only on $\lambda_{1},\lambda_{2}$, and $m$, such that if $\varepsilon_{n}$ is as in (4) satisfying (A.5) and Assumption 2.1 is satisfied, then for all $j=1,\ldots,d$,
$\mathinner{\!\bigl\lvert\mathbb{E}\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{1,j})-\mu_{j}\bigr\rvert}\vee\mathinner{\!\bigl\lvert\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1})-\mu_{j}\bigr\rvert}\leq C\sigma_{m}\varepsilon_{n}^{1-\frac{1}{m}}$ (G.9)
Proof.
Fix $j\in\mathinner{\{1,\ldots,d\}}$ and recall the definitions of $\underline{\alpha}_{j},\overline{\alpha}_{j},\underline{\beta}_{j}$ and $\overline{\beta}_{j}$ from (A.3) and (A.4). Since $\mathbb{E}\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{1,j})-\mu_{j}\leq\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})-\mu_{j}$, (G.9) follows from i) (C.5) of Lemma C.5 in Kock and Preinerstorfer (2025) applied to $X_{1,j}$ with $s_{1}=c_{1,n}\varepsilon_{n}$$s_{2}=1-c_{2,n}\varepsilon_{n}$ and ii) (C.6) of Lemma C.5 in Kock and Preinerstorfer (2025) applied to $X_{1,j}$ with $s_{1}=c_{2,n}\varepsilon_{n}$$s_{2}=1-c_{1,n}\varepsilon_{n}$, recalling that $\sigma_{m,j}\leq\sigma_{m}$ (noting that our Assumption 2.1 implies the assumptions there, and in either case the requirement $s_{1}<s_{2}$ imposed in Lemma C.5 in Kock and Preinerstorfer (2025) follows from (A.5) and Lemma G.2), such that
$\mathinner{\!\bigl\lvert\mathbb{E}\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{1,j})-\mu_{j}\bigr\rvert}\vee\mathinner{\!\bigl\lvert\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})-\mu_{j}\bigr\rvert}$
$\leq 2\sigma_{m}(c_{1,n}\varepsilon_{n})^{1-\frac{1}{m}}+\sigma_{m}\mathinner{\Bigl(1+\mathinner{\Bigl[\frac{c_{2,n}\varepsilon_{n}}{1-c_{2,n}\varepsilon_{n}}\Bigr]}^{\frac{1}{m}}\Bigr)}(c_{2,n}\varepsilon_{n})^{1-\frac{1}{m}}$
$\leq 2\sigma_{m}(c_{1,n}\varepsilon_{n})^{1-\frac{1}{m}}+2\sigma_{m}(c_{2,n}\varepsilon_{n})^{1-\frac{1}{m}},$
the second inequality following from $c_{2,n}\varepsilon_{n}<1/2$ by Lemma G.2 and (A.5) and $(0,1)\ni z\mapsto z/(1-z)$ being increasing. The upper bounds on $c_{1,n}$ and $c_{2,n}$ from Lemma G.2 now yield the conclusion of the lemma. ∎
The following result is similar to Lemma G.2 and bounds $c_{1,n}^{\prime}$ and $c_{2,n}^{\prime}$ as defined in (D.1) and (D.2), respectively. It will be used in Lemma (G.7) below, as well as in the proof of Theorem 3.1.
Lemma G.6.
For $c_{1,n}^{\prime}$ as defined in (D.1), $c_{2,n}^{\prime}$ as defined in (D.2), and $\varepsilon_{n}^{\prime}$ as defined in (7) and satisfying (9), it holds that
$0<(1-\lambda_{1}^{\prime-1})\exp\mathinner{\Bigl({-\frac{1}{\lambda_{2}^{\prime}(1-\lambda_{1}^{\prime-1})}-1}\Bigr)}\leq\$$c_{1,n}^{\prime}<1,$
$1<\$$c_{2,n}^{\prime}\leq 2+\lambda_{2}^{\prime-1}+\sqrt{\lambda_{2}^{\prime-2}+4\lambda_{2}^{\prime-1}},$
and
$0<\varepsilon_{n}^{\prime}\min(c_{1,n}^{\prime},c_{2,n}^{\prime})$ $\leq\varepsilon_{n}^{\prime}(c_{1,n}^{\prime}+c_{2,n}^{\prime})$
$\leq 2\varepsilon_{n}^{\prime}+\frac{\log(d^{2}n)}{n}+\sqrt{\mathinner{\Bigl(\frac{\log(d^{2}n)}{n}\Bigr)}^{2}+4\frac{\log(d^{2}n)}{n}\varepsilon_{n}^{\prime}}.$ (G.10)
Proof.
The proof is almost identical to the one of Lemma G.2, but we now apply Lemma B.3 in Kock and Preinerstorfer (2025) with $\delta=\frac{6}{d^{2}n}$ and $\epsilon=\varepsilon_{n}^{\prime}$. ∎
The following lemma, which follows from Lemma B.5 in Kock and Preinerstorfer (2025), is an analogue to Lemma G.3, replacing $\varepsilon_{n}$ in (4) with $\varepsilon_{n}^{\prime}$ in (7).
Lemma G.7.
Let $Z_{1},\ldots,Z_{n}$ be i.i.d. real random variables and suppose that the random variables $\tilde{Z}_{1},\ldots,\tilde{Z}_{n}$ satisfy
$\mathinner{\!\bigl\lvert\mathinner{\bigl\{i\in\mathinner{\{1,\ldots,n\}}\mathrel{\mathop{\ordinarycolon}}\tilde{Z}_{i}\neq Z_{i}\bigr\}}\bigr\rvert}\leq\overline{\eta}_{n}n.$
Suppose that $\varepsilon_{n}^{\prime}$ is as defined in (7) and satisfies (9). Then, for $c_{1,n}^{\prime}$ as in (D.1) and $c_{2,n}^{\prime}$ as in (D.2), we have
$0<\varepsilon_{n}^{\prime}\min(c_{1,n}^{\prime},c_{2,n}^{\prime})\leq\varepsilon_{n}^{\prime}(c_{1,n}^{\prime}+c_{2,n}^{\prime})<1,$
and each of (G.11)–(G.14) below holds with probability at least $1-\frac{1}{d^{2}n}$:
$\tilde{Z}_{\lceil\varepsilon^{\prime}_{n}n\rceil}^{*}$ $\geq$ $Q_{c_{1,n}^{\prime}\varepsilon^{\prime}_{n}}(Z_{1});$ (G.11)
$\tilde{Z}_{\lceil(1-\varepsilon^{\prime}_{n})n\rceil}^{*}$ $\geq$ $Q_{1-c_{2,n}^{\prime}\varepsilon^{\prime}_{n}}(Z_{1});$ (G.12)
$\tilde{Z}_{\lfloor\varepsilon^{\prime}_{n}n\rfloor+1}^{*}$ $\leq$ $Q_{c_{2,n}^{\prime}\varepsilon^{\prime}_{n}}(Z_{1});$ (G.13)
$\tilde{Z}_{\lfloor(1-\varepsilon^{\prime}_{n})n\rfloor+1}^{*}$ $\leq$ $Q_{1-c_{1,n}^{\prime}\varepsilon^{\prime}_{n}}(Z_{1}).$ (G.14)
Proof.
The first statement in the present lemma follows from (9) along with (G.6) of Lemma G.6. For the remaining statements, apply Lemma B.5 in Kock and Preinerstorfer (2025) with $\delta=\frac{6}{d^{2}n}$ (which is strictly less than one since we assume throughout that $d\geq 2$ and $n\geq 4$, cf. the sentence before Theorem 2.1), $\epsilon=\varepsilon_{n}^{\prime}$, and $\eta$ there being $\overline{\eta}_{n}$. To this end, note that
  1. 1.
    $c_{1}$, and $c_{2}$ there equal $c^{\prime}_{1,n}$ and $c^{\prime}_{2,n}$, respectively, for $\delta=\frac{6}{d^{2}n}$,
  2. 2.
    that $\varepsilon^{\prime}_{n}\geq\lambda^{\prime}_{1}\overline{\eta}_{n}$ by definition and $\varepsilon^{\prime}_{n}c^{\prime}_{2,n}<1$ (which has already been verified),
  3. 3.
    $\varepsilon^{\prime}_{n}<1/2$ follows from (9).